### v3

# Globally-Constrained Decentralized Optimization with Variable Coupling

Abstract
Many realistic decision-making problems in networked scenarios, such as formation control and collaborative task offloading, often involve complicatedly entangled local decisions, which, however, have not been sufficiently investigated yet. Motivated by this, we study a class of decentralized optimization problems with a variable coupling structure that is new to the literature. Specifically, we consider a network of nodes collaborating to minimize a global objective subject to a collection of global inequality and equality constraints, which are formed by the local objective and constraint functions of the nodes. On top of that, we allow such local functions of each node to depend on not only its own decision variable but the decisions of its neighbors as well. To address this problem, we propose a decentralized projected primal-dual algorithm. It first incorporates a virtual-queue technique with a primal-dual-primal scheme, and then linearizes the non-separable objective and constraint functions to enable decentralized implementation. Under mild conditions, we derive $O(1/k)$ convergence rates for both objective error and constraint violations. Finally, two numerical experiments corroborate our theoretical results and illustrate the competitive performance of the proposed algorithm.
Index Terms:
Decentralized optimization, globally-coupled constraint, primal-dual algorithm, variable coupling.
I Introduction
Decentralized optimization has garnered considerable recent attention due to its broad applications in various networked systems such as communication networks[1], smart grids[2], and computing networks[3]. In these networked systems, each node possesses a collection of local data, typically inaccessible to other nodes for privacy protection, and all the nodes aim to make optimal decisions by solving a global optimization problem determined by all their local data. Decentralized optimization techniques allow the nodes to collaboratively solve such a problem through communicating with their neighboring nodes only, leading to high scalability with respect to the network size and the data volume.
To date, a host of decentralized optimization algorithms have been developed. Among these works, the majority focus on consensus optimization, which aims at finding a common decision that minimizes the sum of the local objectives under, mostly, no constraints [4, 5] or rather simple constraints such as set constraints [6, 7], local equality/inequality constraints[8, 9], and equality/inequality constraints known to all the nodes [10, 11]. Another line of research explores problems with local variables subject to coupling equality/inequality constraints [12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25]. In particular, to solve the challenging problems with globally-coupled nonlinear inequality constraints consisting of all the local constraint functions from the nodes, the distributed primal-dual subgradient algorithms [15, 16, 23, 22, 20, 24, 19], the distributed method based on operator splitting[18], the distributed integrated primal-dual proximal algorithm [25], and the relaxation and successive distributed decomposition method[17] along with its extension [21] are developed. These works also derive various convergence results, including asymptotic convergence [15, 16, 17, 18, 19] as well as $O(\ln k/\sqrt{k})$ [20], $O(1/\ln k)$ [21], $O(1/\sqrt{k})$ [22, 23, 24], and $O(1/k)$ [25] convergence rates.
A common limitation of the aforementioned decentralized optimization methods [4, 5, 6, 7, 12, 13, 8, 9, 10, 11, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25] is that each node’s local objective and constraint functions only rely on its own local decision variable. This excludes many realistic problems where each node’s local function values are influenced by the decisions made by other nodes in the network, especially the node’s neighbors. A representative example is the task offloading problem over cooperative mobile edge computing networks [26], where each edge server’s task completion time and local resource consumption are affected by the task offloaded from the neighboring servers. Such a variable coupling structure also arises in collaborative resource allocation [27], distributed model predictive control [28], wireless localization [29], etc.
Motivated by this, we consider a decentralized optimization problem with globally-coupled nonlinear inequality and linear equality constraints. Moreover, for each node, we allow both the local objective function and the local constraint functions to involve variable coupling, i.e., these local functions are associated with the decision variables of the node itself and all its neighbors. This coupling structure endows the optimization problem with a relatively general form that generalizes the prior prevalent decentralized optimization models in [4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25], which do not admit variable coupling in the local functions, and the formulations of the application problems in [26, 27, 28], which admit variable coupling but involve no globally-coupled constraints. A few existing works [30, 31, 32, 33, 34, 35] also consider variable coupling, but they either focus on simpler problems (e.g., [30] on unconstrained problems and [31] on convex feasibility problems) or different variable coupling models [32, 33, 34, 35] that cannot generalize ours (cf. Section II).
In this paper, we develop a decentralized projected primal-dual algorithm for addressing the aforementioned globally-constrained optimization problem with variable coupling. To this end, we first construct a variant of the augmented Lagrangian function with respect to an equivalent form of the original problem, whose gradient has a certain degree of separability. Then, the primal update is designed so as to cause a descent of the augmented Lagrangian variant, where the gradient projection operation is employed. The dual updates can be viewed as a non-trivial integration of ideas from the virtual queue method [36] and a dual variant of the P-EXTRA algorithm [37]. Eventually, the variable coupling structure and the globally-coupled constraints are effectively decoupled, so that our proposed algorithm can be executed in a fully decentralized fashion. We show that the proposed algorithm achieves $O(1/k)$ convergence rates in terms of both objective error and constraint violations under mild conditions. Its effectiveness and efficiency are demonstrated by two numerical examples.
The contributions of this paper are highlighted as follows:
  • To the best of our knowledge, this is the first exploration of decentralized optimization with both variable coupling and globally-coupled constraints.
  • Compared to the existing works [15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25] that are able to solve problems with global inequality constraints but no variable coupling (i.e., special cases of our problem), the $O(1/k)$ convergence rate of our proposed algorithm is stronger than the asymptotic convergence established in [15, 16, 17, 18, 19], faster than the convergence rates in [20, 22, 21, 23, 24], and comparable to the rate in [25].
  • Our algorithm has lower computational costs per iteration than the algorithms in [25, 24, 20, 19, 21, 17, 16, 22] because it does not need to solve an optimization subproblem per iteration. Additionally, our algorithm adopts constant step-sizes, which usually leads to faster convergence in practice compared to the algorithms with diminishing step-sizes [15, 16, 17, 20, 21, 22, 23, 24].
This paper is organized as follows. Section II formulates the problem and presents examples of applications. Section III develops the decentralized projected primal-dual algorithm and Section IV provides the convergence analysis, followed by two numerical experiments in Section V. Finally, Section VI concludes the paper.
[…]
For any differential function $f:\mathbb{R}^{n}\rightarrow\mathbb{R}^{m}$, $\operatorname{dom}(f):=\{x:~{}f(x)<+\infty\}$ is the domain of $f$; $\partial f(x)$ denotes the Jacobian matrix of function $f$ at $x$ and $\nabla f$ is the gradient of $f$ if $m=1$. A function $f$ is said to be $L_{f}$-smooth over a set $X$ if there exists $L_{f}\geq 0$ such that $\|\nabla f(x)-\nabla f(y)\|\leq L_{f}\|x-y\|$, $\forall x,y\in X$. Given $L\geq 0$, $f$ is said to be Lipschitz continuous with Lipschitz constant $L$ over a set $X$ if $\|f(x)-f(y)\|\leq L\|x-y\|$, $\forall x,y\in X$.
II Problem Formulation
This section describes a decentralized globally-constrained optimization problem with variable coupling and presents a couple of application examples.
Consider a network modeled as an undirected, connected graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where the vertex set $\mathcal{V}=\{1,\ldots,n\}$ represents the set of $n$ nodes and the edge set $\mathcal{E}\subseteq\{\{i,j\}:i,j\in\mathcal{V},\;i\neq j\}$ represents the set of bidirectional links. Let $\mathcal{N}_{i}=\{j\in\mathcal{V}:\{i,j\}\in\mathcal{E}~{}or~{}i=j\}$ be the neighbor set of node $i$ including itself.
Suppose all the nodes collaborate to solve the following global convex optimization problem:
$\begin{array}[]{cl}{\underset{x_{i}\in\mathbb{R}^{d_{i}},\;\forall i\in\mathcal{V}}{\operatorname{minimize}}{}}&{\sum\limits_{i\in\mathcal{V}}f_{i}\left(x_{\mathcal{N}_{i}}\right)}\\ {\operatorname{subject~{}to}}&{\sum\limits_{i\in\mathcal{V}}g_{i}\left(x_{\mathcal{N}_{i}}\right)\leq\mathbf{0}_{p}},\\ {}\hfil&{\sum\limits_{i\in\mathcal{V}}}A_{i}x_{\mathcal{N}_{i}}={\sum\limits_{i\in\mathcal{V}}}b_{i},\\ {}\hfil&{x_{i}\in X_{i},\;\forall i\in\mathcal{V}.}\end{array}$ (1)
For each $i\in\mathcal{V}$, $x_{i}\in\mathbb{R}^{d_{i}}$ is the decision variable of node $i$ and is restricted to node $i$’s constraint set $X_{i}\subseteq\mathbb{R}^{d_{i}}$. Additionally, $f_{i}:\mathbb{R}^{\sum_{j\in\mathcal{N}_{i}}d_{j}}\rightarrow\mathbb{R}$ is the local objective function and $g_{i}:\mathbb{R}^{\sum_{j\in\mathcal{N}_{i}}d_{j}}\rightarrow\mathbb{R}^{p}$ is the local inequality constraint function of node $i$, which are known only to node $i$ but are associated with the decision variables of both node $i$ and its neighbors, denoted by $x_{\mathcal{N}_{i}}$. Here, each $g_{i}$ is a vector-valued function, and we express it as $g_{i}\left(x_{\mathcal{N}_{i}}\right)=[g_{i1}\left(x_{\mathcal{N}_{i}}\right),\ldots,g_{ip}\left(x_{\mathcal{N}_{i}}\right)]^{T}$. Moreover, $A_{i}\in\mathbb{R}^{m\times\sum_{j\in\mathcal{N}_{i}}d_{j}}$ and $b_{i}\in\mathbb{R}^{m}$, privately owned by node $i$, encode node $i$’s local component of the global equality constraint in (1), which, again, involves the decision variables of all the nodes in $\mathcal{N}_{i}$.
Compared to most distributed optimization models[4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 30, 31, 32, 33, 34, 35] in the literature, problem (1) has a more complicated and general coupling structure and, thus, is more challenging to solve. Specifically, the problems considered in many existing works [4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25] are “separable” in the sense that each local function belonging to node $i$ is only associated with node $i$’s decision variable $x_{i}$. In contrast, problem (1) is non-separable, as each $f_{i}$, $g_{i}$, and $A_{i}$ are associated with $x_{\mathcal{N}_{i}}$, i.e., $x_{j}$ $\forall j\in\mathcal{N}_{i}$. In addition, $g_{i}$, $A_{i}$, $b_{i}$ $\forall i\in\mathcal{V}$ constitute the globally-coupled inequality and equality constraints in an additive manner. Such a non-separable structure is not uncommon in networked applications, as each node’s cost, utility, capacity, etc., are often influenced by its neighbors’ states and play a role in the network-wide problems. Similar variable coupling structures are also considered in [30, 31], but [30] only investigates unconstrained convex optimization and [31] studies a convex feasibility problem.
Remark 1.
There have been several other coupling models for distributed convex optimization with an additive global objective function. In [32] and [33], each element in a global decision vector is assigned to a subset of nodes that are not necessarily neighbors, and the local objective of each node $i$ is a function of all the elements assigned to node $i$. In [34], each node’s local objective involves not only its own decision variable but also a common aggregative term associated with the decision variables of all the nodes. The global objective in [35] consists of a separable function and a known-to-all function whose variable is a linear combination of all nodes’ decision variables. The above-mentioned objective functions are different from that of problem (1), while problem (1) admits a more general constraint model than those of [32, 33, 34, 35]. In particular, [32, 34, 35] only consider unconstrained problems, [33] only allows for local constraints, yet (1) involves global constraints with coupled variables.
In fact, by letting each node maintain local copies of the decision variables of all the other nodes, problem (1) can be converted into a globally-coupled consensus optimization problem, so that a handful of existing algorithms are applicable (e.g.,[38, 39]). However, such a reformulation would significantly increase the costs in memory, computations, and communications [30], so we do not adopt this approach.
We impose the following assumption on problem (1).
Assumption 1.
Problem (1) satisfies the following conditions:
  1. (a)
    For each $i\in\mathcal{V}$, $X_{i}$ is a compact convex set.
  2. (b)
    For each $i\in\mathcal{V}$, $\sum_{i\in\mathcal{V}}f_{i}\left(x_{\mathcal{N}_{i}}\right)$ and $g_{i}$ are convex on the Cartesian products $X_{\mathcal{V}}$ and $X_{\mathcal{N}_{i}}$, respectively.
  3. (c)
    For each $i\in\mathcal{V}$, $f_{i}$ is $L_{f}$-smooth on $X_{\mathcal{N}_{i}}$ for some $L_{f}\geq 0$.
  4. (d)
    For each $i\in\mathcal{V}$, $g_{ij}$ $\forall j\in\{1,\ldots,p\}$ are $L_{g}$-smooth on $X_{\mathcal{N}_{i}}$ for some $L_{g}\geq 0$.
  5. (e)
    For each $i\in\mathcal{V}$, $g_{i}$ is Lipschitz continuous on $X_{\mathcal{N}_{i}}$ with Lipschitz constant $\beta>0$.
  6. (f)
    There exists an optimum $\mathbf{x}^{\star}=[(x_{1}^{\star})^{T},\ldots,(x_{n}^{\star})^{T}]^{T}\in X_{\mathcal{V}}$ to problem (1).
  7. (g)
    There exist $\tilde{x}_{i}\in\operatorname{rel\;int}(X_{i})$ $\forall i\in\mathcal{V}$ such that ${\sum_{i\in\mathcal{V}}g_{i}\left(\tilde{x}_{\mathcal{N}_{i}}\right)<\mathbf{0}_{p}}$ and ${\sum_{i\in\mathcal{V}}}A_{i}\tilde{x}_{\mathcal{N}_{i}}={\sum_{i\in\mathcal{V}}}b_{i}$.
Assumption 1 indicates that problem (1) is a smooth convex optimization problem with zero duality gap due to the Slater’s condition in Assumption 1(g). Note that we only assume convexity for the global objective function, so that each local objective $f_{i}$ is not necessarily convex. Also, the Lipschitz conditions of each $\nabla f_{i}$, $\nabla g_{ij}$, and $g_{i}$ in Assumption 1(c), 1(d), and 1(e) only need to hold on the compact constraint set $X_{\mathcal{N}_{i}}$. The conditions in Assumption 1 are mild and, later in Section IV-B, we will compare them with the assumptions imposed in the related works.
II-A Examples of Applications
Many real-world engineering problems can be cast into the form of (1). Below, we provide two of such examples.
II-A1 A Multi-Vehicle Formation Stabilization Problem
Consider a multi-vehicle formation stabilization problem. Our goal is to stabilize $n$ vehicles on a plane toward an equilibrium point in a decentralized way. For each vehicle $i$, $s_{i}^{k}=[(p_{i}^{k})^{T},(v_{i}^{k})^{T}]^{T}\in\mathbb{R}^{4}$ is the state and $u_{i}^{k}\in\mathbb{R}^{2}$ is the control input at time $k\geq 0$, where $p_{i}^{k}\in\mathbb{R}^{2}$, $v_{i}^{k}\in\mathbb{R}^{2}$ are the position and the velocity of vehicle $i$, respectively. The system dynamic of each vehicle $i$ is given by
$s_{i}^{k+1}=A_{i}s_{i}^{k}+B_{i}u_{i}^{k}.$ (2)
Given the initial states $s_{i}^{0}$ $\forall i=1,\ldots,n$, all the vehicles obtain their control actions through model predictive control (MPC), which requires them to cooperatively solve a constrained convex optimization problem as follows.
Let $N\geq 1$ be the prediction horizon in MPC. The optimization variable associated with each vehicle $i$ is $x_{i}=[(s_{i}^{1})^{T},\ldots,(s_{i}^{N})^{T},(u_{i}^{0})^{T},\ldots,(u_{i}^{N-1})^{T}]^{T}\in\mathbb{R}^{6N}$. Then, the multi-vehicle formation stabilization problem is formulated as the following optimization problem
$\begin{array}[]{cl}{\underset{x_{i},\;\forall i=1,\ldots,n}{\operatorname{minimize}}{}}&\sum\limits_{i=1}^{n}\Big{(}x_{\mathcal{N}_{i}}^{T}H_{i}x_{\mathcal{N}_{i}}+Q_{i}^{T}x_{\mathcal{N}_{i}}\Big{)}\\ \operatorname{subject~{}to}&\sum_{i=1}^{n}P_{i}x_{i}\leq b_{i},\\ &C_{i}x_{i}=b_{i}^{\prime},\;\forall i=1,\ldots,n,\\ &x_{i}\in X_{i},\;\forall i=1,\ldots,n.\end{array}$ (3)
In (3), each $H_{i}$ is a symmetric positive semidefinite matrix and each $Q_{i}\in\mathbb{R}^{6|\mathcal{N}_{i}|N}$, which characterize the cost of vehicle $i$. The objective in (3) is the sum of tracking cost of each vehicle, where each vehicle $i$’s tracking cost is influenced by the positions of its neighboring vehicles [28, 40]. The first constraint in (3) stands for resource limitations and position/velocity requirements across the whole system. The second constraint represents the dynamics of each vehicle $i$ over the prediction horizon, which is directly derived from (2). The last constraint restricts each $x_{i}$ to some compact convex set $X_{i}$ due to certain system limitations. In (3), $H_{i}$, $Q_{i}$, $P_{i}$, $b_{i}$, $C_{i}$, $b_{i}^{\prime}$, and $X_{i}$ are only accessible to vehicle $i$. Clearly, the distributed MPC problem (3) is a particular form of (1).
II-A2 Task Offloading over Cooperative Mobile Edge Computing Networks
Consider a task offloading problem among $n$ edge servers, which form a connected network. Each edge server receives a large quantity of latency-sensitive tasks from exogenous sources, such as mobile devices. Since the computation resources of each edge server are limited and often distinct, each edge server needs to offload part of the received tasks to neighboring edge servers in order to make full use of computation resources[26]. All the edge servers cooperate to find optimal task offloading decisions by minimizing the total completion time of all the tasks subject to the limitations of channel bandwidth and computation capability.
We label the set of edge servers as $\mathcal{V}=\{1,2,\ldots,n\}$. Let $\mathcal{N}_{i}$ be the neighbor set of edge server $i$ including itself. We define $x_{ii}$ and $x_{ij}$ as the quantity of tasks allocated by edge server $i$ to itself and to its neighboring edge server $j$, respectively. Let $x_{ij}$ $\forall j\in\mathcal{N}_{i}$ constitute the decision vector $x_{i}$ of edge server $i$. Also, let the local objective of edge server $i$ be the task completion time for handling its offloaded task, which, apparently, is determined by the tasks offloaded by both edge server $i$ itself and its neighbors, i.e., $x_{ji}$ $\forall j\in\mathcal{N}_{i}$. Thus, we can write the local objective of edge server $i$ as $f_{i}\left(x_{\mathcal{N}_{i}}\right)$ for simplicity. The coupled inequality constraints $\sum_{i\in\mathcal{V}}g_{i}\left(x_{\mathcal{N}_{i}}\right)\leq\mathbf{0}_{p}$ are often associated with the total channel bandwidth or computation resource limit across the network, where $g_{i}$ represents the edge server $i$’s channel bandwidth or computation resource consumption [41]. For each edge server, the sum of the task offloaded to itself and neighboring servers should be equal to the tasks it received from the exogenous sources, which can be described by linear equality constraints. Therefore, such a task offloading problem is in the form of (1).
III Algorithm Development
This section develops a decentralized algorithm to solve problem (1).
III-A Problem Transformation
We first transfer (1) into an equivalent form to facilitate our algorithm design.
By adding $n$ auxiliary variables $t_{i}\in\mathbb{R}^{p}$ $\forall i\in\mathcal{V}$, the inequality constraint $\sum_{i\in\mathcal{V}}g_{i}(x_{\mathcal{N}_{i}})\leq\mathbf{0}_{p}$ can be converted to $g_{i}(x_{\mathcal{N}_{i}})\leq t_{i}$ $\forall i\in\mathcal{V}$ and $\sum_{i\in\mathcal{V}}t_{i}=\mathbf{0}_{p}$. Let
$\mathbf{y}$ $=[y_{1}^{T},\ldots,y_{n}^{T}]^{T}\in\mathbb{R}^{N},$ (4)
$y_{i}$ $=[x_{i}^{T},t_{i}^{T}]^{T},\forall i\in\mathcal{V},$ (5)
where $N=np+\sum_{i\in\mathcal{V}}d_{i}$. Then, problem (1) can be equivalently transformed into
$\begin{array}[]{cl}\underset{\mathbf{y}\in\mathbb{R}^{N}}{\operatorname{minimize}}{}&\mathbf{f}(\mathbf{y}):=\sum\limits_{i\in\mathcal{V}}f_{i}(x_{\mathcal{N}_{i}})\\ \operatorname{subject~{}to}&\mathbf{G}(\mathbf{y}):=\left[\begin{array}[]{c}g_{1}(x_{\mathcal{N}_{1}})\!-\!t_{1}\\ \vdots\\ g_{n}(x_{\mathcal{N}_{n}})\!-\!t_{n}\end{array}\right]\!\leq\!\mathbf{0}_{np},\\ &(\mathbf{1}_{n}\otimes I_{m+p})^{T}(\mathbf{B}\mathbf{y}-\mathbf{c})=\mathbf{0}_{m+p},\\ &\mathbf{y}\in Y,\end{array}$ (6)
where $Y$ is the Cartesian product of $Y_{i}=\{[x_{i}^{T},t_{i}^{T}]^{T}:x_{i}\in X_{i},\ t_{i}\in\mathbb{R}^{p}\}$ $\forall i\in\mathcal{V}$, $\mathbf{B}=\operatorname{diag}(B_{1},\ldots,B_{n})$ with $B_{i}=\operatorname{diag}(\sum_{j\in\mathcal{N}_{i}}A_{ji},I_{p})$, and $\mathbf{c}=[c_{1}^{T},\ldots,c_{n}^{T}]^{T}$ with $c_{i}=[b_{i}^{T},\mathbf{0}_{p}^{T}]^{T}$. Here, we partition each $A_{i}$ into $A_{ij}\in\mathbb{R}^{m\times d_{j}}$ $\forall j\in\mathcal{N}_{i}$ with each $A_{ij}$ formed by $d_{j}$ columns of $A_{i}$.
Assumption 1 on problem (1) leads to the following properties of the equivalent problem (6).
Proposition 1.
Suppose Assumption 1 holds. Then, problem (6) satisfies the following properties:
  1. (a)
    $\mathbf{f}(\mathbf{y})$ and $\mathbf{G}(\mathbf{y})$ are convex on the set $Y$.
  2. (b)
    $\nabla\mathbf{f}$ is Lipschitz continuous with Lipschitz constant $L_{F}:=L_{f}\max_{i\in\mathcal{V}}(|\mathcal{N}_{i}|)$ on the set $Y$.
  3. (c)
    $\mathbf{G}(\mathbf{y})$ is Lipschitz continuous with Lipschitz constant $\tilde{\beta}:=\sqrt{(1+\beta^{2})(\max_{i\in\mathcal{V}}(|\mathcal{N}_{i}|))}$ on the set $Y$.
  4. (d)
    There exists at least one optimum $\mathbf{y}^{\star}$ to problem (6).
  5. (e)
    Strong duality holds between problem (6) and its Lagrange dual problem.
Proof.
See Appendix -A. ∎
III-B Algorithm Design
This subsection designs an algorithm for solving the equivalent problem (6).
Our proposed algorithm maintains four variables $\mathbf{y}^{k}\in\mathbb{R}^{N}$, $\mathbf{q}^{k}\in\mathbb{R}^{np}$, $\mathbf{u}^{k}\in\mathbb{R}^{n(m+p)}$, and $\mathbf{z}^{k}\in\mathbb{R}^{n(m+p)}$, where $\mathbf{y}^{k}$ is the primal variable, $\mathbf{q}^{k}$ and $\mathbf{u}^{k}$ are the dual variables associated with the inequality constraint and the equality constraint in (6), respectively, and $\mathbf{z}^{k}$ is the auxiliary variable that enables the update of $\mathbf{u}^{k}$.
The algorithm parameters include two constant step-sizes $\gamma,\rho>0$ and two weight matrices $W,H$ that facilitate information fusion across the network and comply with the network topology. We let $W=P^{W}\otimes I_{m+p},~{}H=P^{H}\otimes I_{m+p}$, and require $P^{W},P^{H}\in\mathbb{R}^{n\times n}$ meet the assumption below.
Assumption 2.
(a) $[P^{W}]_{ij}=[P^{H}]_{ij}=0$, $\forall i\in\mathcal{V}$, $\forall j\notin\mathcal{N}_{i}$; (b) $P^{W}$ and $P^{H}$ are symmetric and positive semidefinite; (c) $P^{W}\mathbf{1}_{n}=\mathbf{1}_{n}$, $\operatorname{Null}(P^{H})=\operatorname{span}(\mathbf{1}_{n})$; (d) $P^{W}+P^{H}\prec I_{n}$.
There are many options for $P^{W},P^{H}$ under Assumption 2. For example, we may let $P^{W}\!=\!\frac{I+P^{\prime}}{2}$ and $P^{H}\!=\!\frac{I-P^{\prime}}{2}$ for some $P^{\prime}\in\mathbb{R}^{n\times n}$ such that $[P^{\prime}]_{ij}=[P^{\prime}]_{ji}>0$$\forall\{i,j\}\in\mathcal{E}$, $\sum_{j\in\mathcal{N}_{i}}[P^{\prime}]_{ij}=1$, $[P^{\prime}]_{ii}>0$, and $[P^{\prime}]_{ij}=0$ otherwise [25].
At initial iteration $k=0$, we arbitrarily select $\mathbf{y}^{0}\in Y$ and $\mathbf{u}^{0}\in\mathbb{R}^{n(m+p)}$. Then, we initialize $\mathbf{z}^{0}=\rho H\mathbf{u}^{0}$ and $\mathbf{q}^{0}=\max\{-\mathbf{G}(\mathbf{y}^{0}),\mathbf{0}_{np}\}$. At each subsequent iteration $k+1$, all the variables are updated according to
$\mathbf{y}^{k+1}$ $=\mathcal{P}_{Y}[\mathbf{y}^{k}-\gamma\cdot(\mathbf{d}^{k})^{T}],$ (7)
$\mathbf{q}^{k+1}$ $=\max\{-\mathbf{G}(\mathbf{y}^{k+1}),\mathbf{q}^{k}\!+\!\mathbf{G}(\mathbf{y}^{k+1})\},$ (8)
$\mathbf{u}^{k+1}$ $=W\mathbf{u}^{k}+\frac{1}{\rho}(\mathbf{B}\mathbf{y}^{k+1}-\mathbf{c}-\mathbf{z}^{k}),$ (9)
$\mathbf{z}^{k+1}$ $=\mathbf{z}^{k}+\rho H\mathbf{u}^{k+1}.$ (10)
In the primal update (7), $\mathbf{d}^{k}$ is the Jacobian matrix (or equivalently, $(\mathbf{d}^{k})^{T}$ is the gradient) of the following augmented-Lagrangian-like function $\mathcal{R}^{k}(\mathbf{y})$ at $\mathbf{y}^{k}$:
$\mathcal{R}^{k}(\mathbf{y})=$ $\mathbf{f}(\mathbf{y})+\langle\mathbf{q}^{k}+\mathbf{G}(\mathbf{y}^{k}),\mathbf{G}(\mathbf{y})\rangle$
$+\langle W\mathbf{u}^{k}-\frac{1}{\rho}\mathbf{z}^{k},\mathbf{B}\mathbf{y}-\mathbf{c}\rangle+\frac{1}{2\rho}\|\mathbf{B}\mathbf{y}-\mathbf{c}\|^{2},$ (11)
and the expression of $\mathbf{d}^{k}$ is thus given by
$\mathbf{d}^{k}=$ $\frac{\partial\mathbf{f}(\mathbf{y}^{k})}{\partial\mathbf{y}}+(W\mathbf{u}^{k}-\frac{1}{\rho}\mathbf{z}^{k})^{T}\mathbf{B}+\frac{1}{\rho}(\mathbf{y}^{k})^{T}\mathbf{B}^{T}\mathbf{B}$
$+(\mathbf{q}^{k}+\mathbf{G}(\mathbf{y}^{k}))^{T}\frac{\partial\mathbf{G}(\mathbf{y}^{k})}{\partial\mathbf{y}}+\frac{1}{\rho}\mathbf{c}^{T}\mathbf{B},$ (12)
where $\frac{\partial\mathbf{f}(\mathbf{y}^{k})}{\partial\mathbf{y}}$ and $\frac{\partial\mathbf{G}(\mathbf{y}^{k})}{\partial\mathbf{y}}$ are the Jacobian matrices of $\mathbf{f}(\mathbf{y})$ and $\mathbf{G}(\mathbf{y})$ at $\mathbf{y}^{k}$, respectively. Clearly, the primal update (7) is a gradient projection operation aiming at the descent of the augmented-Lagrangian-like function $\mathcal{R}^{k}(\mathbf{y})$ in (11).
The dual update (8) adopts the idea of virtual queue in [36] (yet our primal update is completely different from that in [36]), which allows $\mathbf{q}^{k}$ to emulate the Lagrange multiplier associated with the inequality constraint $\mathbf{G}(\mathbf{y})\leq\mathbf{0}$ in (6) and benefits the convergence rate analysis compared to conventional dual subgradient updates.
To understand the rationale behind the remaining dual updates (9) and (10), we consider the dual problem of (6) in absence of the inequality constraint $\mathbf{G}(\mathbf{y})\leq\mathbf{0}$ [^1] The Lagrange dual problem is originally given by $\operatorname{maximize}_{u\in\mathbb{R}^{m+p}}\min_{\mathbf{y}\in\mathbb{R}^{N}}\mathbf{f}(\mathbf{y})+\langle u,(\mathbf{1}_{n}\otimes I_{m+p})^{T}(\mathbf{B}\mathbf{y}-\mathbf{c})\rangle$. Here we consider its equivalent form (13) for our algorithm design.:
$\begin{split}\underset{\mathbf{u}\in\mathbb{R}^{n(m+p)}}{\operatorname{maximize}}~{}~{}&~{}D_{e}(\mathbf{u}):=\min_{\mathbf{y}\in Y}\mathbf{f}(\mathbf{y})+\langle\mathbf{u},\mathbf{B}\mathbf{y}-\mathbf{c}\rangle\\ \operatorname{subject~{}to}~{}&~{}~{}u_{1}=\ldots=u_{n},\end{split}$ (13)
where $\mathbf{u}=[u_{1}^{T},\ldots,u_{n}^{T}]^{T}$. We first make an attempt to solve the dual problem (13) by employing the distributed first-order method P-EXTRA [42]. According to [37], such a dual P-EXTRA algorithm gives
$\mathbf{u}^{k+1}=$ $\operatorname{\arg\;\min}_{\mathbf{u}\in\mathbb{R}^{n(m+p)}}-D_{e}(\mathbf{u})+\langle\mathbf{u},\mathbf{z}^{k}\rangle$
$+\frac{\rho}{2}\|\mathbf{u}-W\mathbf{u}^{k}\|^{2},$ (14)
where $\mathbf{z}^{k}$ takes the same recursive form as (10) with $\mathbf{z}^{0}=\rho H\mathbf{u}^{0}$. Note that the above updates are not implementable because in general the dual function $D_{e}(\mathbf{u})$ is inaccessible.
To overcome this issue, we propose the following primal-dual-primal approach: First, note that we may rewrite (14) as
$\mathbf{z}^{k}+\rho(\mathbf{u}^{k+1}-W\mathbf{u}^{k})\in\partial D_{e}(\mathbf{u}^{k+1}).$
On the other hand, if we set
$\mathbf{y}^{k+1}\in\operatorname{\arg\;\min}_{\mathbf{y}\in Y}\mathbf{f}(\mathbf{y})+\langle\mathbf{u}^{k+1},\mathbf{B}\mathbf{y}-\mathbf{c}\rangle,$ (15)
then $\mathbf{B}\mathbf{y}^{k+1}-\mathbf{c}$ is also a subgradient of $D_{e}(\mathbf{u}^{k+1})$ like $\mathbf{z}^{k}+\rho(\mathbf{u}^{k+1}-W\mathbf{u}^{k})$. Hence, we choose to equalize these two items, leading to the update equation of $\mathbf{u}^{k+1}$ which is exactly (9). Then, to derive a legitimate update equation for $\mathbf{y}^{k}$, note that (15) can be rewritten as $-\mathbf{B}^{T}\mathbf{u}^{k+1}\in\partial\mathbf{f}(\mathbf{y}^{k+1})$, which, along with (9), results in $-\mathbf{B}^{T}(W\mathbf{u}^{k}+\frac{1}{\rho}(\mathbf{B}\mathbf{y}^{k+1}-\mathbf{c}-\mathbf{z}^{k}))\in\partial\mathbf{f}(\mathbf{y}^{k+1})$. This is equivalent to
$\mathbf{y}^{k+1}\in\operatorname{\arg\;\min}_{\mathbf{y}\in Y}\mathcal{R}^{k}(\mathbf{y}),$ (16)
where $\mathcal{R}^{k}(\mathbf{y})$ is given in (11) (with $\mathbf{G}(\mathbf{y})\equiv\mathbf{0}$). Therefore, (16), (9), and (10) together constitute an equivalent form of the aforementioned dual P-EXTRA algorithm, i.e., (14) and (10). Recall that (7) is a gradient projection operation for emulating (16). Consequently, for problem (6) with equality constraints only, our proposed algorithm can be viewed as a dual first-order method that originates from dual P-EXTRA yet introduces the gradient projection operation to acquire readily available and computationally efficient estimates of dual subgradients.
Remark 2.
Our proposed algorithm does not directly minimize $\mathcal{R}^{k}(\mathbf{y})$, i.e., adopt (16) as the primal update. This is because $f_{i}$ and $g_{i}$ appearing in $\mathcal{R}^{k}(\mathbf{y})$ are functions of coupled variables, which disable the decentralized realization of (16). Similarly, the primal update of IPLUX in [25] minimizes $\mathcal{R}^{k}(\mathbf{y})$ with $\mathbf{f}(\mathbf{y})$ substituted by its first-order approximation, which can only tackle problems without variable coupling and, thus, is unable to solve problem (1) without any centralized coordination. In contrast, the gradient of $\mathcal{R}^{k}(\mathbf{y})$ has a decomposable structure, so that the gradient projection step (7) can be executed decentralizedly, which will be elaborated in Section III-C. Also, (7) can significantly reduce the computational complexity of (16). Nevertheless, by adopting (7) instead of (16), we indeed compromise on the accuracy in estimating dual subgradients, thereby intensifying the challenge of convergence analysis.
III-C Decentralized Implementation
Below we describe the decentralized implementation of the proposed algorithm (7)–(10).
Let each node $i\in\mathcal{V}$ maintain $y_{i}^{k}\in\mathbb{R}^{d_{i}+p},q_{i}^{k}\in\mathbb{R}^{p},u_{i}^{k}\in\mathbb{R}^{m+p}$, and $z_{i}^{k}\in\mathbb{R}^{m+p}$, where $y_{i}^{k},q_{i}^{k},u_{i}^{k},z_{i}^{k}$ are the $i$-th block of the variables $\mathbf{y}^{k},\mathbf{q}^{k},\mathbf{u}^{k},\mathbf{z}^{k}$, respectively. We further partition $y_{i}^{k}=[(x_{i}^{k})^{T},(t_{i}^{k})^{T}]^{T}$, where $x_{i}^{k}\in\mathbb{R}^{d_{i}}$ and $t_{i}^{k}\in\mathbb{R}^{p}$.
To obtain the update of $y_{i}^{k}$, let $\mathbf{d}^{k}=[d_{1}^{k},\ldots,d_{n}^{k}]$, where $d_{i}^{k}=\frac{\partial\mathcal{R}^{k}(\mathbf{y}^{k})}{\partial y_{i}}\!\in\!\mathbb{R}^{1\times(d_{i}+p)}$ $\forall i\in\mathcal{V}$ represents the Jacobian matrix of $\mathcal{R}^{k}(\mathbf{y})$ with respect to $y_{i}^{k}$. The expression of $d_{i}^{k}$ $\forall i\in\mathcal{V}$ directly follows from (12) and needs to compute $\frac{\partial\mathcal{R}^{k}(\mathbf{y}^{k})}{\partial y_{i}}$ by utilizing only elementary matrix operations, which is given
$\begin{split}d_{i}^{k}&=\Big{(}(\sum\limits_{j\in\mathcal{N}_{i}}\![P^{W}]_{ij}u_{j}^{k})^{T}-\frac{1}{\rho}(z_{i}^{k})^{T}\Big{)}B_{i}+[\tilde{d}_{i1}^{k},~{}\tilde{d}_{i2}^{k}],\end{split}$ (17)
where
$\tilde{d}_{i1}^{k}$ $=\sum\limits_{j\in\mathcal{N}_{i}}\!\!\Big{(}\frac{\partial f_{j}(x^{k}_{\mathcal{N}_{j}})}{\partial x_{i}}+(q_{j}^{k}+g_{j}(x_{\mathcal{N}_{j}}^{k})-t_{j}^{k})^{T}\!\frac{\partial g_{j}(x^{k}_{\mathcal{N}_{j}})}{\partial x_{i}}\Big{)}$
$~{}~{}~{}~{}~{}+\frac{1}{\rho}\big{(}\!\sum\limits_{j\in\mathcal{N}_{i}}\!\!\!A_{ji}x_{i}^{k}\big{)}^{T}\sum\limits_{j\in\mathcal{N}_{i}}A_{ji}+\frac{1}{\rho}b_{i}^{T}\sum\limits_{j\in\mathcal{N}_{i}}A_{ji},$ (18)
$\begin{split}\tilde{d}_{i2}^{k}&=\frac{1}{\rho}(t_{i}^{k})^{T}-(q_{i}^{k}+g_{i}(x_{\mathcal{N}_{i}}^{k})-t_{i}^{k})^{T}.\end{split}$ (19)
The above equations (17)–(19) indicate that $d_{i}^{k}$ can be obtained by information interaction between neighboring nodes. Oracle 1 presents the decentralized calculation of $d_{i}^{k}$ $\forall i\in\mathcal{V}$.
By utilizing the structure of the projection operation, $\mathbf{y}^{k+1}$ in (7) is decomposed as the following local updates:
$y_{i}^{k+1}=\mathcal{P}_{Y_{i}}[y_{i}^{k}-\gamma\cdot(d_{i}^{k})^{T}]~{}~{}\forall i\!\in\!\mathcal{V}.$ (20)
Owing to the neighbor sparse structures of $W,H$ and the definition of $\mathbf{G}$ in (6), the decentralized updates of $\mathbf{q}^{k+1}$, $\mathbf{u}^{k+1}$, and $\mathbf{z}^{k+1}$ in (8)–(10) can be easily rewritten as the following local operations: For each $k\geq 0$ and each $i\in\mathcal{V}$,
$q_{i}^{k+1}$ $=\max\left\{t_{i}^{k+1}\!-g_{i}(x_{\mathcal{N}_{i}}^{k+1}\!),~{}q_{i}^{k}\!+\!g_{i}(x_{\mathcal{N}_{i}}^{k+1}\!)\!-\!t_{i}^{k+1}\right\},$ (21)
$u_{i}^{k+1}$ $=\sum_{j\in\mathcal{N}_{i}}[P^{W}]_{ij}u_{j}^{k}+\frac{1}{\rho}\left(B_{i}y_{i}^{k+1}-c_{i}-z_{i}^{k}\right),$ (22)
$z_{i}^{k+1}$ $=z_{i}^{k}+\rho\sum_{j\in\mathcal{N}_{i}}[P^{H}]_{ij}u_{j}^{k+1}.$ (23)
Algorithm 1 describes the detailed implementation of (20)–(23) taken by all the nodes. Given arbitrary $y_{i}^{0}\in Y_{i}$, $u_{i}^{0}\in\mathbb{R}^{p}$ $\forall i\in\mathcal{V}$, each node $i\!\in\!\mathcal{V}$ calculates $z_{i}^{0},q_{i}^{0}$ by collecting $u_{j}^{0},x_{j}^{0}$ from its every neighbor $j\in\mathcal{N}_{i}$. Then, at each $k\geq 0$, each node $i\in\!\!\mathcal{V}$ updates its primal variable $y_{i}^{k+1}$ according to (20), which requires node $i$ to calculate $d_{i}^{k}$ first. According to Oracle 1, each node $i$ needs to collect $x_{j}^{k}$, $u_{j}^{k}$, $A_{ji}$, and $\frac{\partial f_{j}(x^{k}_{\mathcal{N}_{j}})}{\partial x_{i}}+(q_{j}^{k}+g_{j}(x_{\mathcal{N}_{j}}^{k})-t_{j}^{k})^{T}\frac{\partial g_{j}(x^{k}_{\mathcal{N}_{j}})}{\partial x_{i}}\!$ from every neighbor node $j$ for obtaining $\tilde{d}_{i1}^{k},\tilde{d}_{i2}^{k}$ and then computes $d_{i}^{k}$ based on $\tilde{d}_{i1}^{k},\tilde{d}_{i2}^{k}$. The updates of $q_{i}^{k+1},u_{i}^{k+1}$ in (21), (22) can be realized by requesting $x_{j}^{k+1}$ from neighbor $j$. In addition, each node $i\in\mathcal{V}$ computes $z_{i}^{k+1}$ according to (23) by collecting $u_{j}^{k+1}$ from its neighbor $j$. Obviously, the updates (20)–(23) can be realized in a decentralized and inexpensive way.
TABLE 1: Oracle 1: Decentralized calculation of $d_{i}^{k}$ $\forall i\in\mathcal{V}$ $\forall k\geq 0$.
Upon receiving $x_{j}^{k}$, $u_{j}^{k}$ and $A_{ji}$ from every neighbor $j$, each node $i\in\mathcal{V}$
(a) calculates and sends $\frac{\partial f_{i}(x^{k}_{\mathcal{N}_{i}})}{\partial x_{j}}+(q_{i}^{k}+g_{i}(x_{\mathcal{N}_{i}}^{k})\!-t_{i}^{k})^{T}\frac{\partial g_{i}(x^{k}_{\mathcal{N}_{i}})}{\partial x_{j}}$ to its neighbor $j$;
(b) computes $\tilde{d}_{i1}^{k}$ according to (18);
(c) computes $\tilde{d}_{i2}^{k}$ according to (19);
(d) computes $d_{i}^{k}$ according to (17).
Algorithm 1 Decentralized Projected Primal-Dual Algorithm
1:  Initialization: Each node $i\in\mathcal{V}$ arbitrary selects $u_{i}^{0}\in\mathbb{R}^{m+p}$ and $y_{i}^{0}=[(x_{i}^{0})^{T},(t_{i}^{0})^{T}]^{T}\in Y_{i}$.
2:  Each node $i\!\in\!\mathcal{V}$ sends $x_{i}^{0},\!\!\ u_{i}^{0}$ and $A_{ij}$ to every neighbor $j$ and sets $z_{i}^{0}\!=\rho\sum_{j\in\mathcal{N}_{i}}[P^{H}]_{ij}u_{j}^{0}$, $q_{i}^{0}=\max\{t_{i}^{0}-g_{i}(x_{\mathcal{N}_{i}}^{0}),\mathbf{0}_{p}\}$.
3:  for $k=0,1,2,\ldots$, each node $i\in\mathcal{V}$ do
4:     Compute $d_{i}^{k}$ according to Oracle 1;
5:     Update $y_{i}^{k+1}$ according to (20);
6:     Send $x_{i}^{k+1}$ to every neighbor $j$;
7:     Update $q_{i}^{k+1}$ according to (21);
8:     Update $u_{i}^{k+1}$ according to (22);
9:     Send $u_{i}^{k+1}$ to every neighbor $j$;
10:     Update $z_{i}^{k+1}$ according to (23).
11:  end for
III-C1 Special case of no variable coupling
It is noteworthy that during initialization and each iteration $k\geq 0$, each node $i\in\mathcal{V}$ needs to communicate with its neighbors in lines 2, 4, 6, 9 of Algorithm 1. Nevertheless, if each $f_{i}$, $g_{i}$ only depends on $x_{i}$ and each $A_{i}$ only involves $x_{i}$, i.e., no variable coupling exists, the calculation of $d_{i}^{k}$ in Oracle 1 requires no information except $u_{j}^{k}$ from each neighbor $j$, which greatly simplifies the calculation of $d_{i}^{k}$. Moreover, in this case, updating $q_{i}^{k+1}$ in line 7 of Algorithm 1 does not rely on $x_{j}^{k}$ received from neighbor $j$, and updating $u_{i}^{k+1}$ in line 8 involves only $A_{i}$, no need of $A_{ji}$. The above simplified operators require each node $i\in\mathcal{V}$ to send only $u_{i}^{k}$ to its every neighbor $j$, and the transmission of $x_{i}^{0},A_{ij}$ in line 2, step (a) in Oracle 1, line 6 of Algorithm 1 can be canceled. Consequently, the communication and computational costs of Algorithm 1 can be significantly reduced. Additionally, in this case, all the nodes execute these simplified updates without disclosing their primal decisions, local objectives, and local constraint functions, thereby highly preserving information privacy.
IV Convergence Analysis
This section is dedicated to analyzing the convergence performance of Algorithm 1, along with a comparison to related works and a discussion on the parameter selections for guaranteeing the theoretical results.
IV-A Convergence Results
To establish the convergence analysis, we define $\mathbf{z}^{\star}=\mathbf{B}\mathbf{y}^{\star}\!-\mathbf{c}$ and let $(\bm{\lambda}^{\star},u^{\star})$ be an optimal solution of the Lagrange dual problem of problem (6), where $\bm{\lambda}^{\star}\geq\mathbf{0}_{np}$. Also, let $\mathbf{u}^{\star}=\mathbf{1}_{n}\otimes u^{\star}$. We keep track of the running average $\bar{\mathbf{y}}^{k}:=\frac{1}{k}\sum\limits_{\ell=1}^{k}\mathbf{y}^{\ell}$ $\forall k\geq 1$ to establish the convergence analysis. The following two theorems state the convergence results of Algorithm 1.
TABLE 1: Table 1: Comparison with related methods in solving problems with globally-coupled inequality constraints yet with no variable coupling in local functions.
Algorithm compact domain smooth problem existence of equality constraint local optimization per iteration constant step-size convergence result
consensus-based primal-dual perturbation method [15] $\surd$ $\surd$ asymptotic
dual decomposition based method [16] $\surd$ $\surd$ asymptotic
RSDD [17] $\surd$ $\surd$ asymptotic
primal-dual gradient method [18] $\surd$ $\surd$ $\surd$ $\surd$ asymptotic
ALT [19] $\surd$ $\surd$ $\surd$ asymptotic
dual subgradient method [20] $\surd$ $\surd$ $\surd$ $O(\ln k/\sqrt{k})$
DPD-TV[21] $\surd$ $\surd$ $O(1/\ln k)$
$\text{DSA}_{2}$ [22] $\surd$ $\surd$ $O(1/\sqrt{k})$
C-SP-SG [23] $\surd$ $O(1/\sqrt{k})$
B-DPP [24] $\surd$ $\surd$ $O(1/\sqrt{k})$
IPLUX [25] $\surd$ $\surd$ $\surd$ $O(1/k)$
Algorithm 1 $\surd$ $\surd$ $\surd$ $\surd$ $O(1/k)$
Theorem 1.
Suppose Assumptions 12 hold. Let $\gamma,\rho>0$ be such that $C:=\frac{1}{2\rho}\|\mathbf{z}^{0}-\mathbf{z}^{\star}\|_{H^{\dagger}}^{2}+\frac{\rho}{2}(\|\mathbf{u}^{0}\|_{W}+\|\mathbf{u}^{\star}\|_{W})^{2}+\frac{1}{2}\|\mathbf{G}(\mathbf{y}^{0})\|^{2}+\frac{1}{2}\|\mathbf{G}(\mathbf{y}^{\star})\|^{2}+\frac{1}{2}\|\mathbf{q}^{0}\|^{2}+\|\bm{\lambda}^{\star}\|^{2}+\frac{1}{2}(\mathbf{y}^{0}-\mathbf{y}^{\star})^{T}(\frac{I}{\gamma}-\frac{\mathbf{B^{T}}\mathbf{B}}{\rho})(\mathbf{y}^{0}-\mathbf{y}^{\star})\geq 0$. Also suppose
$\frac{1}{\gamma}\geq\frac{1}{\rho}\|\mathbf{B}^{T}\mathbf{B}\|+\tilde{\beta}^{2}+L_{F}+4\sqrt{np}(\|\bm{\lambda}^{\star}\|+\sqrt{C})L_{g},$ (24)
where $L_{F}$, $\tilde{\beta}$, and $L_{g}$ are given in Proposition 1(b)(c) and Assumption 1(d), respectively. Then,
$\mathbf{G}(\bar{\mathbf{y}}^{k})\leq\frac{2(\|\bm{\lambda}^{\star}\|+\sqrt{C})}{k}\mathbf{1}_{np},\quad\forall k\geq 1,$ (25)
$\|(\mathbf{1}_{n}\otimes I_{m+p})^{T}(\mathbf{B}\bar{\mathbf{y}}^{k}-\mathbf{c})\|\leq\frac{D^{0}}{k},\quad\forall k\geq 1.$ (26)
where $D^{0}=\rho\sqrt{n}(\|\mathbf{u}^{0}\|_{W}+\|\mathbf{u}^{\star}\|_{W}+\sqrt{2C/\rho})$.
Proof.
See Appendix -B. ∎
The existence of $\gamma,\rho>0$ under the parameter conditions in Theorem 1 is guaranteed and will be justified in Section IV-C. Additionally, although those theoretical parameter conditions involve the unavailable primal and dual optima, Section IV-C demonstrates that they can be satisfied based on accessible information only.
Theorem 2.
Under all the conditions in Theorem 1,
$-\frac{C^{0}}{k}\leq\mathbf{f}(\bar{\mathbf{y}}^{k})\!-\!\mathbf{f}(\mathbf{y}^{\star})\leq\frac{S^{0}}{k},\quad\forall k\geq 1,$ (27)
where
$C^{0}=$ $2(\|\bm{\lambda}^{\star}\|+\sqrt{C})\mathbf{1}_{np}^{T}\bm{\lambda}^{\star}$
$+\rho\|\mathbf{u}^{\star}\|(\|\mathbf{u}^{0}\|_{W}+\|\mathbf{u}^{\star}\|_{W}+\sqrt{2C/\rho}),$
$S^{0}=$ $\frac{1}{2\rho}\|\mathbf{z}^{0}-\mathbf{z}^{\star}\|_{H^{\dagger}}^{2}+\frac{\rho}{2}\|\mathbf{u}^{0}\|_{W}^{2}+\frac{1}{2}\|\mathbf{y}^{0}-\mathbf{y}^{\star}\|^{2}_{\frac{I}{\gamma}-\frac{\mathbf{B^{T}}\mathbf{B}}{\rho}}$
$+\frac{1}{2}\|\mathbf{q}^{0}\|^{2}-\frac{1}{2}\|\mathbf{G}(\mathbf{y}^{0})\|^{2}.$
Proof.
See Appendix -C. ∎
Theorem 1 and Theorem 2 show that the proposed algorithm achieves $O(1/k)$ convergence rates for both constraint violations and objective function value to problem (6). Due to the equivalence between problem (1) and problem (6), the above results also indicate $O(1/k)$ convergence of constraint violations and objective function value to problem (1), which is described in Corollary 1. In Corollary 1, $\bar{x}_{\mathcal{N}_{i}}^{k}$ $\forall k\geq 1$ is the vector obtained by stacking $\bar{x}_{j}^{k}:=\frac{1}{k}\sum_{\ell=1}^{k}x_{j}^{\ell}$ $\forall j\in\mathcal{N}_{i}$.
Corollary 1.
Suppose all the conditions in Theorem 1 hold. For each $k\geq 1$,
$\sum_{i\in\mathcal{V}}g_{i}(\bar{x}_{\mathcal{N}_{i}}^{k})\leq\frac{2n(\|\bm{\lambda}^{\star}\|+\sqrt{C})+D^{0}}{k}\mathbf{1}_{p},$ (28)
$\|{\sum_{i\in\mathcal{V}}}(A_{i}\bar{x}_{\mathcal{N}_{i}}^{k}-b_{i})\|\leq\frac{D^{0}}{k},$ (29)
$-\frac{C^{0}}{k}\leq\Bigl{(}\sum_{i\in\mathcal{V}}f_{i}(\bar{x}_{\mathcal{N}_{i}}^{k})\Bigr{)}-f^{\star}\leq\frac{S^{0}}{k},$ (30)
where $C$, $D^{0}$, $C^{0}$, and $S^{0}$ are defined in Theorems 1 and 2, and $f^{\star}$ is the optimal value of problem (1).
Proof.
See Appendix -D. ∎
IV-B Comparative Discussion
To the best of our knowledge, no prior methods can be directly applied to solve problem (1) in a decentralized way. Nevertheless, the algorithms in [15, 16, 17, 18, 20, 22, 21, 19, 23, 24, 25] are able to solve problems that admit globally-coupled inequality constraints yet no variable coupling, where each node’s local objective and constraint functions only depend on its own decision variable and are independent of its neighbors’ decisions. Here, we compare Algorithm 1 with those algorithms for solving such specialized forms of (1).
Observe from Table 1 that the algorithms in [15, 16, 17, 18, 19] do not have guaranteed convergence rates. The convergence rate results in [20, 22, 23, 21, 24] are weaker than the $O(1/k)$ rate for Algorithm 1, while they allow certain nonsmoothness in their problems. Although IPLUX [25] achieves $O(1/k)$ convergence rates in no need of a compact domain, it needs every node to solve a local convex optimization problem at each iteration (same for the algorithms in [24, 20, 19, 21, 17, 16, 22]), so that it has much higher computational complexity than Algorithm 1. Notably, only Algorithm 1 and [18, 20, 25, 19] consider problems with both inequality and equality constraints, while the others do not include equality constraints in their problem formulations. Although we may transform each equality constraint into two inequalities with reverse signs, this does not work for [15, 23, 21] because their methods require each inequality constraint to be strictly negative at a Slater point in order to bound the dual optimal set, and would significantly increase the variable dimensions and computational/communication complexities of the algorithms in [16, 17, 22, 24]. Furthermore, Algorithm 1 and [19, 18, 25] adopt constant step-sizes, whereas the remaining algorithms in Table 1 use diminishing step-sizes, which may lead to slow convergence in practice.
IV-C Parameter Selections
In this subsection, we show that there exist $\gamma,\rho>0$ satisfying the parameter conditions in Theorem 1, i.e., $C\geq 0$ and (24), and then discuss how to select the parameters under those theoretical conditions in practice.
For simplicity, we arbitrarily fix $\rho>0$. To achieve $C\geq 0$, we may simply let $0<\gamma\leq\frac{\rho}{\|\mathbf{B}^{T}\mathbf{B}\|}$. When all the inequality constraint functions $g_{i}(x_{\mathcal{N}_{i}})$ $\forall i\in\mathcal{V}$ are affine, (24) reduces to $0<\gamma\leq(\tilde{\beta}^{2}+L_{F}+\frac{1}{\rho}\|\mathbf{B}^{T}\mathbf{B}\|)^{-1}$ because $L_{g}=0$. When $g_{i}(x_{\mathcal{N}_{i}})$ $\forall i\in\mathcal{V}$ are nonlinear, below we provide a range for $\gamma$ that sufficiently guarantees (24).
To do so, define $D:=\tilde{\beta}^{2}+L_{F}+4\sqrt{np}L_{g}\|\bm{\lambda}^{\star}\|+\frac{1}{\rho}\|\mathbf{B}^{T}\mathbf{B}\|$ and $R:=C+\frac{1}{2\rho}\|\mathbf{y}^{0}-\mathbf{y}^{\star}\|_{\mathbf{B^{T}}\mathbf{B}}^{2}-\frac{1}{2\gamma}\|\mathbf{y}^{0}-\mathbf{y}^{\star}\|^{2}\geq 0$ for convenience. Note from the definition of $C$ in Theorem 1 that $D$ and $R$ are both independent of $\gamma$. Then, (24) can be rewritten as
$\gamma D+4\gamma\sqrt{np}L_{g}\sqrt{C}\leq 1.$ (31)
Next, we find an upper bound on the left-hand side of (31). Note that $\gamma D+4\gamma\sqrt{np}\sqrt{C}L_{g}\leq\gamma D+4\sqrt{np}L_{g}\sqrt{\gamma^{2}R+\frac{\gamma}{2}\|\mathbf{y}^{0}-\mathbf{y}^{\star}\|^{2}}\leq\gamma(D+4\sqrt{np}L_{g}\sqrt{R})+4\sqrt{\gamma}\sqrt{np}L_{g}\|\mathbf{y}^{0}-\mathbf{y}^{\star}\|$. Clearly, setting this upper bound no more than $1$ suffices to guarantee (31) and equivalently (24), which can be achieved by letting
$0<\gamma\leq\tilde{\gamma},$ (32)
where $\tilde{\gamma}:=\Bigl{(}\sqrt{4L_{g}^{2}np\|\mathbf{y}^{0}-\mathbf{y}^{\star}\|^{2}+(D+4\sqrt{np}L_{g}\sqrt{R})}+2L_{g}\sqrt{np}\|\mathbf{y}^{0}-\mathbf{y}^{\star}\|\Bigr{)}^{-2}$ is independent of $\gamma$ and $\tilde{\gamma}\leq\frac{\rho}{\|\mathbf{B}^{T}\mathbf{B}\|}$. Therefore, for any given $\rho>0$, there always exists a sufficiently small $\gamma\in(0,\tilde{\gamma}]$ such that $C\geq 0$ and (24) holds.
Such a sufficiently small $\gamma$ can be obtained either empirically (e.g., by trial and error) or analytically (e.g., by estimating $\tilde{\gamma}$). To do the latter, we may construct some lower bound on $\tilde{\gamma}$ without the knowledge of the optima $\mathbf{y}^{\star},\mathbf{u}^{\star},\bm{\lambda}^{\star},\mathbf{z}^{\star}$, leading to a subinterval of $(0,\tilde{\gamma}]$ that still ensures $C\geq 0$ and (24). The main challenge here is to bound $\|\mathbf{y}^{\star}\|$, $\|\mathbf{u}^{\star}\|_{W}$, $\|\bm{\lambda}^{\star}\|$, $\|\mathbf{z}^{\star}\|$ by means of available information only, and some feasible approaches are provided as follows.
  1. (1)
    Upper bound of $\|\mathbf{y}^{\star}\|$: Recall that $\mathbf{y}^{\star}$ is composed of $\mathbf{x}^{\star}$ and $t^{\star}$. We may utilize the diameter of the compact set $X_{\mathcal{V}}$ and any point in $X_{\mathcal{V}}$ to bound $\|\mathbf{x}^{\star}\|$. Additionally, we may choose $t_{i}^{\star}=g_{i}(x_{\mathcal{N}_{i}}^{\star})-\frac{1}{n}\sum_{j\in\mathcal{V}}g_{j}(x_{\mathcal{N}_{j}}^{\star})$ $\forall i\in\mathcal{V}$ (see proof of Proposition 1(d) for its optimality and feasibility), and its upper bound can be acquired from the continuity of each $g_{j}$ over the compact $X_{\mathcal{N}_{j}}$.
  2. (2)
    Upper bound of $\|\mathbf{z}^{\star}\|$: Since $\mathbf{z}^{\star}=\mathbf{B}\mathbf{y}^{\star}\!-\mathbf{c}$, its upper bound can be directly obtained using the bound on $\|\mathbf{y}^{\star}\|$.
  3. (3)
    Upper bound of $\|\bm{\lambda}^{\star}\|$: By virtue of the Slater’s condition of problem (6) and the continuity of $\sum_{i\in\mathcal{V}}f_{i}(x_{\mathcal{N}_{i}})$ over the compact $X_{\mathcal{V}}$, the method in [43] can be employed to derive an upper bound on $\|\bm{\lambda}^{\star}\|$.
  4. (4)
    Upper bound of $\|\mathbf{u}^{\star}\|_{W}$: Note that the Lagrangian function of (6) can be expressed as $L(\mathbf{y},\bm{\lambda},\mathbf{u})=\mathbf{f}(\mathbf{y})+\langle\bm{\lambda},\mathbf{G}(\mathbf{y})\rangle+\langle\mathbf{u},\mathbf{B}\mathbf{y}-\mathbf{c}\rangle$, where $\mathbf{u}=[u_{1}^{T},\ldots,u_{n}^{T}]^{T}$ with $u_{1}=\cdots=u_{n}$. Let the dual optima be written as $\bm{\lambda}^{\star}=[(\bm{\lambda}_{1}^{\star})^{T},\cdots,(\bm{\lambda}_{n}^{\star})^{T}]^{T}\in\mathbb{R}^{np}$ and $u_{i}^{\star}=[(u_{i1}^{\star})^{T},(u_{i2}^{\star})^{T}]^{T}$ $\forall i\in\mathcal{V}$ with $u_{i1}^{\star}\in\mathbb{R}^{m}$ and $u_{i2}^{\star}\in\mathbb{R}^{p}$. From the stationarity condition in the Karush-Kuhn-Tucker Conditions, we obtain $u_{i2}^{\star}=\bm{\lambda}_{i}^{\star}$ $\forall i\in\mathcal{V}$ and
    $\nabla_{\mathbf{x}}\mathbf{f}(\mathbf{y}^{\star})+\nabla_{\mathbf{x}}\mathbf{G}(\mathbf{y}^{\star})\bm{\lambda}^{\star}+\bar{A}\bar{\mathbf{u}}_{1}^{\star}=\mathbf{0},$ (33)
    where $\bar{A}=\operatorname{diag}(\bar{A}_{1},\cdots,\bar{A}_{n})$ with each $\bar{A}_{i}=(\sum_{j\in\mathcal{N}_{i}}A_{ji})^{T}$ and $\bar{\mathbf{u}}_{1}^{\star}=[(u_{11}^{\star})^{T},\cdots,(u_{n1}^{\star})^{T}]^{T}$. Since $W\mathbf{u}^{\star}=\mathbf{u}^{\star}$, we have $\|\mathbf{u}^{\star}\|^{2}_{W}=\|\mathbf{u}^{\star}\|^{2}=\sum_{i=1}^{n}(\|u_{i1}^{\star}\|^{2}+\|u_{i2}^{\star}\|^{2})=\|\bar{\mathbf{u}}_{1}^{\star}\|^{2}+\|\bm{\lambda}^{\star}\|^{2}$. Hence, it remains to bound $\|\bar{\mathbf{u}}_{1}^{\star}\|$. This may be realized by viewing (33) as a linear equation with $\bar{\mathbf{u}}_{1}^{\star}$ being the unknown, which is guaranteed to have a solution, and deriving the expression of any solution. For example, when each $\bar{A}_{i}$ has full column rank, which occurs in many real-world problems such as task offloading over mobile edge computing networks [26] and economic dispatch for power systems [44], the linear equation (33) has a unique solution
    $\bar{\mathbf{u}}_{1}^{\star}=-(\bar{A}^{T}\bar{A})^{-1}\bar{A}^{T}(\nabla_{\mathbf{x}}\mathbf{f}(\mathbf{y}^{\star})+\nabla_{\mathbf{x}}\mathbf{G}(\mathbf{y}^{\star})\bm{\lambda}^{\star}).$
    When each $\bar{A}_{i}$ has full row rank, the minimum-norm solution is given by
    $\bar{\mathbf{u}}_{1}^{\star}=-\bar{A}^{T}(\bar{A}\bar{A}^{T})^{-1}\!(\nabla_{\mathbf{x}}\mathbf{f}(\mathbf{y}^{\star})+\nabla_{\mathbf{x}}\mathbf{G}(\mathbf{y}^{\star})\bm{\lambda}^{\star}).$
    Thus, using the aforementioned bounds on $\mathbf{y}^{\star},\bm{\lambda}^{\star}$ and the smoothness of $\mathbf{f},\mathbf{G}$, we can find an upper bound on $\|\bar{\mathbf{u}}_{1}^{\star}\|$ and therefore an upper bound on $\|\mathbf{u}^{\star}\|^{2}_{W}=\|\mathbf{u}^{\star}\|^{2}$.
Upon bounding the above primal and dual optima, we are able to estimate the remaining quantities in $\tilde{\gamma}$, which involve certain problem and graph characteristics as well as the initial states, through mild global coordination or proper distributed computation methods. Such additional costs are minor in comparison with solving the complicated optimization problem (1).
V Numerical Experiments
In this section, we validate our theoretical results by applying Algorithm 1 to two numerical examples. We first consider a globally-coupled constrained problem with no variable coupling, which allows us to compare the proposed algorithm with related works. The second numerical experiment intends to verify the effectiveness of Algorithm 1 on handling both variable coupling and globally-coupled constraints.
Figure 1: Network topology with $n=50$ nodes
V-A First Numerical Example
Consider the following problem
$\begin{array}[]{cl}{\underset{x_{i}\in[0,1]}{\operatorname{minimize}}{}}&{\sum\limits_{i\in\mathcal{V}}}c_{i}x_{i}\\ {\operatorname{subject~{}to}}&{\sum\limits_{i\in\mathcal{V}}}-d_{i}\operatorname{log}(1+x_{i})\leq-b.\end{array}$ (34)
Problem (34) arises in, for example, the quality service of wireless networks and the plug-in electric vehicles charging problem [22, 23]. In this simulation, we consider a undirected network with $n=50$ nodes (see Figure 1), and the set of nodes is $\mathcal{V}=\{1,\ldots,50\}$. For each $i\in\mathcal{V}$, the constants $c_{i},d_{i}$ are chosen randomly from a uniform distribution in $[0,1]$ and $b=5$. Clearly, the linear objective and smooth constraints in problem (34) meet Assumption 1(a)–(e). The Slater’s condition can be satisfied when $x_{i}=1~{}\forall i\in\mathcal{V}$. Problem (34) satisfies not only Assumption 1 but also the assumptions for the dual subgradient method [20], $\text{DSA}_{2}$ [22] and C-SP-SG [23].
(a) Optimal error at $\mathbf{x}^{k}$
(b) Optimal error at $\bar{\mathbf{x}}^{k}$
Figure 2: Convergence performance of Algorithm 1 and three alternative methods for solving problem (34).
Based on above settings, we compare Algorithm 1 with the dual subgradient method [20], $\text{DSA}_{2}$ [22] and C-SP-SG [23] since they all have convergence rate guarantees for solving problem (34). For a fair comparison, we hand-tune all the algorithm parameters to make the algorithms achieve the best possible convergence performance. The optimal solution of problem (34) $\mathbf{x}^{\star}=[(x_{1}^{\star})^{T},\ldots,(x_{n}^{\star})^{T}]^{T}$ is obtained by using the solver fmincon in Matlab. Figure 2 plots the optimal errors $\|\mathbf{x}^{k}-\mathbf{x}^{\star}\|$ and $\|\bar{\mathbf{x}}^{k}-\mathbf{x}^{\star}\|$, respectively, where $\mathbf{x}^{k}=[(x_{1}^{k})^{T},(x_{2}^{k})^{T},\ldots,(x_{n}^{k})^{T}]^{T}$ is the network-wide primal iterate at iteration $k$, and $\bar{\mathbf{x}}^{k}:=\frac{1}{k}\sum\limits_{\ell=1}^{k}\mathbf{x}^{\ell}$ is its running average. From Figure 2, we observe that Algorithm 1 converges faster than the other three algorithms, especially for the optimal error $\|\mathbf{x}^{k}-\mathbf{x}^{\star}\|$ at the primal iterate, demonstrating the superior convergence performance of our proposed algorithm.
V-B Second Numerical Example
Consider the following problem
$\begin{array}[]{cl}{\underset{x_{i}\in\mathbb{R}^{d},\forall i\in\mathcal{V}}{\operatorname{minimize}}{}}&\!\sum\limits_{i\in\mathcal{V}}(x_{\mathcal{N}_{i}})^{T}P_{i}x_{\mathcal{N}_{i}}+Q_{i}^{T}x_{\mathcal{N}_{i}}\\ {\operatorname{subject~{}to}}&\!\sum\limits_{i\in\mathcal{V}}(x_{\mathcal{N}_{i}})^{T}\!\!A_{i}x_{\mathcal{N}_{i}}\!+\!a_{i}^{T}x_{\mathcal{N}_{i}}\!\leq\!\sum\limits_{i\in\mathcal{V}}\!d_{i},\\ {}\hfil&\!\sum\limits_{i\in\mathcal{V}}B_{i}x_{\mathcal{N}_{i}}={\sum\limits_{i\in\mathcal{V}}}b_{i},\\ {}\hfil&\!x_{i}\in\{x\in\mathbb{R}^{d}:l_{i}\leq x\leq u_{i}\},\end{array}$ (35)
where each $P_{i},A_{i}\in\mathbb{R}^{d|\mathcal{N}_{i}|\times d|\mathcal{N}_{i}|}$ is symmetric positive semidefinite for guaranteeing the convexity of local functions. Besides, $Q_{i},a_{i}\in\mathbb{R}^{d|\mathcal{N}_{i}|}$, $B_{i}\in\mathbb{R}^{m\times d|\mathcal{N}_{i}|},b_{i}\in\mathbb{R}^{m}$, and $d_{i}\in\mathbb{R}$. These problem data are randomly generated under the above conditions. Under the same network environment as Figure 1, let $m=d=2$ and $X_{i}=\{x\in\mathbb{R}^{d}:l_{i}\leq x\leq u_{i}\}$, $f_{i}\left(x_{\mathcal{N}_{i}}\right)=(x_{\mathcal{N}_{i}})^{T}P_{i}x_{\mathcal{N}_{i}}+Q_{i}^{T}x_{\mathcal{N}_{i}}$ and $g_{i}\left(x_{\mathcal{N}_{i}}\right)=(x_{\mathcal{N}_{i}})^{T}A_{i}x_{\mathcal{N}_{i}}+a_{i}^{T}x_{\mathcal{N}_{i}}-d_{i}$. Apparently, problem (35) satisfies Assumption 1 under proper problem data.
Figure 3: Convergence performance of Algorithm 1 for solving problem (35).
Problem (35) has a complicated coupling structure in the form of (1) and no existing decentralized optimization algorithms can solve (35). Hence, we only execute Algorithm 1 to solve it. Figure 3 depicts the optimal errors at $\mathbf{x}^{k}$ and $\bar{\mathbf{x}}^{k}$, respectively. The results in Figure 3 verify the effectiveness of Algorithm 1 on solving problems with both variable coupling and globally-coupled constraints.
VI Conclusion
In this paper, we have developed a decentralized projected primal-dual algorithm for addressing a generally-coupled constrained optimization problem with a variable coupling structure, where the local objective and constraint functions of each node are determined by local variables as well as the variables of its neighbors. The proposed algorithm is developed based on the gradient projection method, a virtual queue technique, and a primal-dual-primal method, and is shown to achieve convergence at $O(1/k)$ rates in terms of both optimality and infeasibility. The effectiveness and efficiency of the proposed algorithm have been demonstrated via two sets of simulations.
-A Proof of Proposition 1
First, the convexity of $\mathbf{f}$ and $\mathbf{G}$ in (a) is a direct consequence of the convexity of $\sum_{i\in\mathcal{V}}f_{i}\left(x_{\mathcal{N}_{i}}\right)$ and $g_{i}$ in Assumption 1(g).
[…]
$\sum_{i\in\mathcal{V}}\|x_{\mathcal{N}_{i}}-x_{\mathcal{N}_{i}}^{\prime}\|^{2}=\sum_{i\in\mathcal{V}}\sum\limits_{j\in\mathcal{N}_{i}}\|x_{j}-x_{j}^{\prime}\|^{2}$
$\leq\max_{i\in\mathcal{V}}(|\mathcal{N}_{i}|)\sum_{i\in\mathcal{V}}\|x_{i}-x_{i}^{\prime}\|^{2}.$ (36)
It follows that $\|\nabla\mathbf{f}(\mathbf{y})-\nabla\mathbf{f}(\mathbf{y^{\prime}})\|^{2}\leq L_{f}^{2}(\max_{i\in\mathcal{V}}(|\mathcal{N}_{i}|))^{2}\sum_{i\in\mathcal{V}}\|x_{i}-x_{i}^{\prime}\|^{2}\leq L_{f}^{2}(\max_{i\in\mathcal{V}}(|\mathcal{N}_{i}|))^{2}\|\mathbf{y}-\mathbf{y^{\prime}}\|^{2}$, i.e., property (b) holds.
From Assumption 1(e) and (36),
$\|\mathbf{G}(\mathbf{y})-\mathbf{G}(\mathbf{y}^{\prime})\|^{2}$
$=$ $\sum_{i\in\mathcal{V}}\|(g_{i}(x_{\mathcal{N}_{i}})-t_{i})-(g_{i}(x_{\mathcal{N}_{i}}^{\prime})-t_{i}^{\prime})\|^{2}$
$\leq$ $\sum_{i\in\mathcal{V}}(1+\frac{1}{\beta^{2}})\|g_{i}(x_{\mathcal{N}_{i}})-g_{i}(x_{\mathcal{N}_{i}}^{\prime})\|^{2}$
$+\sum_{i\in\mathcal{V}}(1+\beta^{2})\|t_{i}-t_{i}^{\prime}\|^{2}$
$\leq$ $\sum_{i\in\mathcal{V}}(1+\beta^{2})(\|x_{\mathcal{N}_{i}}-x_{\mathcal{N}_{i}}^{\prime}\|^{2}+\|t_{i}-t_{i}^{\prime}\|^{2})$
$\leq$ $(1+\beta^{2})\max_{i\in\mathcal{V}}(|\mathcal{N}_{i}|)\|\mathbf{y}-\mathbf{y^{\prime}}\|^{2},$
where the second step comes from $-2\langle g_{i}(x_{\mathcal{N}_{i}})-g_{i}(x_{\mathcal{N}_{i}}^{\prime}),t_{i}-t_{i}^{\prime}\rangle\leq\frac{1}{\beta^{2}}\|g_{i}(x_{\mathcal{N}_{i}})-g_{i}(x_{\mathcal{N}_{i}}^{\prime})\|^{2}+\beta^{2}\|t_{i}-t_{i}^{\prime}\|^{2}$ $\forall i\in\mathcal{V}$. Therefore, property (c) holds.
To prove property (d), let $\mathbf{x}^{\star}=[(x_{1}^{\star})^{T},\ldots,(x_{n}^{\star})^{T}]$ be an optimum of problem (1), which is guaranteed to exist due to Assumption 1(f). Then, we let $t_{i}^{\star}=g_{i}(x_{\mathcal{N}_{i}}^{\star})-\frac{1}{n}\sum_{j\in\mathcal{V}}g_{j}(x_{\mathcal{N}_{j}}^{\star})$ $\forall i\in\mathcal{V}$, and let $\mathbf{y}^{\star}=[(x_{1}^{\star})^{T},(t_{1}^{\star})^{T},\ldots,(x_{n}^{\star})^{T},(t_{n}^{\star})^{T}]^{T}$. It can be shown that $\mathbf{y}^{\star}$ is feasible to problem (6). Moreover, because $\mathbf{f}(\mathbf{y}^{\star})=\sum_{i\in\mathcal{V}}f_{i}(x_{\mathcal{N}_{i}}^{\star})$ and because (1) and (6) share the same optimal value, $\mathbf{y}^{\star}$ is an optimum of problem (6).
Finally, let $\tilde{x}_{i}$ $\forall i\in\mathcal{V}$ be given by Assumption 1(g), and let $\tilde{t}_{i}=g_{i}(\tilde{x}_{\mathcal{N}_{i}})-\frac{1}{n}\sum_{j\in\mathcal{V}}g_{j}(\tilde{x}_{\mathcal{N}_{j}})$ $\forall i\in\mathcal{V}$. It can thus be verified that $[(\tilde{x}_{1})^{T},(\tilde{t}_{1})^{T},\ldots,(\tilde{x}_{n})^{T},(\tilde{t}_{n})^{T}]^{T}$ is a Slater’s point for problem (6), so that property (e) holds.
-B Proof of Theorem 1
We first bound the constraint violations with respect to problem (6) at $\bar{\mathbf{y}}^{k}$ in the lemma below.
Lemma 1.
Suppose Assumptions 1 and 2 hold. Then, for each $k\geq 1$,
$\mathbf{G}(\bar{\mathbf{y}}^{k})$ $\leq\frac{\mathbf{q}^{k}}{k},$ (37)
$\!\!\!(\mathbf{1}_{n}\otimes I_{m+p})^{T}(\mathbf{B}\bar{\mathbf{y}}^{k}-\mathbf{c})$ $=\frac{\rho}{k}(\mathbf{1}_{n}\otimes I_{m+p})^{T}(\mathbf{u}^{k}-\mathbf{u}^{0}).$ (38)
Proof.
Refer to the proof of [25, eq.(37), eq.(38)]. ∎
From Lemma 1, Theorem 1 can be proved as long as we can further bound $\mathbf{q}^{k}$ and $(\mathbf{1}_{n}\otimes I_{m+p})^{T}(\mathbf{u}^{k}-\mathbf{u}^{0})$. To this end, we introduce the following lemmas.
Lemma 2.
Suppose Assumptions 1 and 2 hold. For each $k\geq 0$, $\nabla\mathcal{R}^{k}(\mathbf{y})$ is Lipschitz continuous on the set $Y$ with Lipschitz constant $L_{\mathcal{R}^{k}}:=L_{F}+\langle\mathbf{q}^{k}+\mathbf{G}(\mathbf{y}^{k}),L_{g}\mathbf{1}_{np}\rangle+\frac{1}{\rho}\|\mathbf{B}^{T}\mathbf{B}\|$.
Proof.
For any $\mathbf{y},\mathbf{y^{\prime}}\in Y$,
$\|\nabla\mathcal{R}^{k}(\mathbf{y})-\nabla\mathcal{R}^{k}(\mathbf{y^{\prime}})\|$
$\leq\|\nabla\mathbf{f}(\mathbf{y})-\nabla\mathbf{f}(\mathbf{y^{\prime}})\|+\|\frac{1}{\rho}\mathbf{B}^{T}\mathbf{B}(\mathbf{y}-\mathbf{y^{\prime}})\|\!$
$+\sum_{i=1}^{n}\sum\limits_{\ell=1}^{p}\|(q_{i\ell}^{k}+G_{i\ell}(\mathbf{y}^{k}))(\nabla G_{i\ell}(\mathbf{y})-\nabla G_{i\ell}(\mathbf{y^{\prime}}))\|,$ (39)
where $q_{i\ell}^{k}$ and $G_{i\ell}$, $\ell=1,\ldots,p$ represent the $\ell$th coordinates of $\mathbf{q}^{k}$ and $G_{i}$, respectively. Note from $\mathbf{q}^{0}=\max\{-\mathbf{G}(\mathbf{y}^{0}),\mathbf{0}\}$ and (8) that $\mathbf{q}^{k}+\mathbf{G}(\mathbf{y}^{k})\geq\mathbf{0}$. It follows from Assumption 1(d) that $\|(q_{i\ell}^{k}+G_{i\ell}(\mathbf{y}^{k}))(\nabla G_{i\ell}(\mathbf{y})-\nabla G_{i\ell}(\mathbf{y^{\prime}}))\|=(q_{i\ell}^{k}+G_{i\ell}(\mathbf{y}^{k}))\|\nabla G_{i\ell}(\mathbf{y})-\nabla G_{i\ell}(\mathbf{y^{\prime}})\|\leq(q_{i\ell}^{k}+G_{i\ell}(\mathbf{y}^{k}))L_{g}\|\mathbf{y}-\mathbf{y^{\prime}}\|$. Incorporating this and Proposition 1(b) into (39) leads to $\|\nabla\mathcal{R}^{k}(\mathbf{y})-\nabla\mathcal{R}^{k}(\mathbf{y^{\prime}})\|\leq L_{\mathcal{R}^{k}}\|\mathbf{y}-\mathbf{y^{\prime}}\|$. ∎
Lemma 3.
For each $k\geq 0$,
$\mathbf{q}^{k}\geq\mathbf{0},$ (40)
$\|\mathbf{q}^{0}\|\leq\|\mathbf{G}(\mathbf{y}^{0})\|,\quad\|\mathbf{q}^{k+1}\|\geq\|\mathbf{G}(\mathbf{y}^{k+1})\|.$ (41)
Proof.
It follows from [36, Lemma 3 and Lemma 4]. ∎
Lemma 4.
Suppose all the conditions in Theorem 1 hold. If there is some $K\geq 1$ such that $\|\mathbf{q}^{k}\|\leq 2(\|\bm{\lambda}^{\star}\|+\sqrt{C})$ $\forall k=0,1,2,\ldots,K-1$, then,
$\sum_{k=1}^{K}(\mathbf{f}(\mathbf{y}^{k})-\mathbf{f}(\mathbf{y}^{\star}))\leq S^{0}-S^{K},$ (42)
where $S^{k}=\frac{1}{2\rho}\|\mathbf{z}^{k}-\mathbf{z}^{\star}\|_{H^{\dagger}}^{2}+\frac{\rho}{2}\|\mathbf{u}^{k}\|_{W}^{2}+\frac{1}{2}\|\mathbf{y}^{k}-\mathbf{y}^{\star}\|^{2}_{\frac{I}{\gamma}-\frac{\mathbf{B^{T}}\mathbf{B}}{\rho}}+\frac{1}{2}\|\mathbf{q}^{k}\|^{2}-\frac{1}{2}\|\mathbf{G}(\mathbf{y}^{k})\|^{2}$ $\forall k\geq 0$.
Proof.
To prove (42), we first bound $\mathbf{f}(\mathbf{y}^{k+1})-\mathbf{f}(\mathbf{y}^{\star})$ $\forall k\geq 0$. Recall from Section III that $(\mathbf{d}^{k})^{T}$ is the gradient of $\mathcal{R}^{k}(\mathbf{y})$ at $\mathbf{y}^{k}$. Thus, (7) can be rewritten as $\mathbf{y}^{k+1}=\operatorname{arg\;min}_{\mathbf{y}\in Y}M^{k}(\mathbf{y})$, where $M^{k}(\mathbf{y})=\mathcal{R}^{k}(\mathbf{y}^{k})+\nabla\mathcal{R}^{k}(\mathbf{y}^{k})^{T}(\mathbf{y}-\mathbf{y}^{k})+\frac{1}{2\gamma}\|\mathbf{y}-\mathbf{y}^{k}\|^{2}$. Since $M^{k}(\mathbf{y})$ is $\frac{1}{\gamma}$-strongly convex and $\mathbf{y}^{k+1}$ minimizes $M^{k}(\mathbf{y})$, we have $M^{k}(\mathbf{y}^{\star})\geq M^{k}(\mathbf{y}^{k+1})+\frac{1}{2\gamma}\|\mathbf{y}^{k+1}-\mathbf{y}^{\star}\|^{2}$, which is equivalent to
$\langle\nabla\mathcal{R}^{k}(\mathbf{y}^{k}),\mathbf{y}^{k+1}-\mathbf{y}^{\star}\rangle$ $\leq\frac{1}{2\gamma}(\|\mathbf{y}^{k}-\mathbf{y}^{\star}\|^{2}-\|\mathbf{y}^{k+1}-\mathbf{y}^{\star}\|^{2})$
$-\frac{1}{2\gamma}\|\mathbf{y}^{k+1}-\mathbf{y}^{k}\|^{2}.$ (43)
On the other hand, from Lemma 2,
$\mathcal{R}^{k}(\mathbf{y}^{k+1})\leq$ $\mathcal{R}^{k}(\mathbf{y}^{k})+\langle\nabla\mathcal{R}^{k}(\mathbf{y}^{k}),\mathbf{y}^{k+1}-\mathbf{y}^{k}\rangle$
$+\frac{L_{\mathcal{R}^{k}}}{2}\|\mathbf{y}^{k+1}-\mathbf{y}^{k}\|^{2}.$ (44)
In addition, due to the convexity of $\mathcal{R}^{k}(\mathbf{y})-\frac{1}{2\rho}\|\mathbf{B}\mathbf{y}-\mathbf{c}\|^{2}$, we obtain
$-\mathcal{R}^{k}(\mathbf{y}^{\star})+\frac{1}{2\rho}\|\mathbf{B}\mathbf{y}^{\star}-\mathbf{c}\|^{2}\leq-\mathcal{R}^{k}(\mathbf{y}^{k})+\frac{1}{2\rho}\|\mathbf{B}\mathbf{y}^{k}-\mathbf{c}\|^{2}$
$+\langle\nabla\mathcal{R}^{k}(\mathbf{y}^{k})-\frac{1}{\rho}\mathbf{B}^{T}(\mathbf{B}\mathbf{y}^{k}-\mathbf{c}),\mathbf{y}^{k}-\mathbf{y}^{\star}\rangle.$
By adding the above inequality and (44), we have
$\mathcal{R}^{k}(\mathbf{y}^{k+1})-\mathcal{R}^{k}(\mathbf{y}^{\star})\leq\langle\nabla\mathcal{R}^{k}(\mathbf{y}^{k}),\mathbf{y}^{k+1}\!-\!\mathbf{y}^{\star}\rangle$
$+\frac{L_{\mathcal{R}^{k}}}{2}\|\mathbf{y}^{k+1}-\mathbf{y}^{k}\|^{2}-\frac{1}{2\rho}\|\mathbf{B}\mathbf{y}^{k}-\mathbf{B}\mathbf{y}^{\star}\|^{2}.$
Let $M_{1}^{k}(\mathbf{y})=\langle W\mathbf{u}^{k}\!-\frac{1}{\rho}\mathbf{z}^{k},\mathbf{B}\mathbf{y}\!-\!\mathbf{c}\rangle+\frac{1}{2\rho}\|\mathbf{B}\mathbf{y}-\mathbf{c}\|^{2}\!$, $M_{2}^{k}(\mathbf{y})=\langle\mathbf{q}^{k}+\mathbf{G}(\mathbf{y}^{k}),\mathbf{G}(\mathbf{y})\rangle$ and we have $\mathcal{R}^{k}(\mathbf{y})=\mathbf{f}(\mathbf{y})+M_{1}^{k}(\mathbf{y})+M_{2}^{k}(\mathbf{y})$. Substituting this definition into the combination of the above inequality and (43) yields
$\mathbf{f}(\mathbf{y}^{k+1})-\!\mathbf{f}(\mathbf{y}^{\star})$
$\leq M_{1}^{k}(\mathbf{y}^{\star})-M_{1}^{k}(\mathbf{y}^{k+1})+M_{2}^{k}(\mathbf{y}^{\star})-M_{2}^{k}(\mathbf{y}^{k+1})$
$+\!(\frac{L_{\mathcal{R}^{k}}}{2}\!-\frac{1}{2\gamma}\!)\|\mathbf{y}^{k+1}\!\!-\!\mathbf{y}^{k}\|^{2}\!-\!\frac{1}{2\rho}\|\mathbf{B}\mathbf{y}^{k}-\mathbf{B}\mathbf{y}^{\star}\|^{2}$
$+\frac{1}{2\gamma}(\|\mathbf{y}^{k}\!-\!\mathbf{y}^{\star}\|^{2}-\|\mathbf{y}^{k+1}\!\!-\mathbf{y}^{\star}\|^{2}).$ (45)
By referring to the proof of [25, Lemma 2], we are able to derive the bounds on $M_{1}^{k}(\mathbf{y}^{\star})-M_{1}^{k}(\mathbf{y}^{k+1})$ and $M_{2}^{k}(\mathbf{y}^{\star})-M_{2}^{k}(\mathbf{y}^{k+1})$ in (45) as follows:
$M_{1}^{k}(\mathbf{y}^{\star})-M_{1}^{k}(\mathbf{y}^{k+1})$
$\leq\frac{1}{2\rho}(\|\mathbf{z}^{\star}-\mathbf{z}^{k}\|_{H^{\dagger}}^{2}-\|\mathbf{z}^{\star}-\mathbf{z}^{k+1}\|_{H^{\dagger}}^{2})$
$+\frac{\rho}{2}(\|\mathbf{u}^{k}\|_{W}^{2}-\|\mathbf{u}^{k+1}\|_{W}^{2})+\frac{1}{2\rho}\|\mathbf{B}\mathbf{y}^{\star}-\mathbf{B}\mathbf{y}^{k+1}\|^{2},$
$M_{2}^{k}(\mathbf{y}^{\star})\!-\!M_{2}^{k}(\mathbf{y}^{k+1})$
$\leq\frac{1}{2}(\|\mathbf{q}^{k}\|^{2}\!-\!\|\mathbf{q}^{k+1}\|^{2})+\!\frac{1}{2}(\|\mathbf{G}(\mathbf{y}^{k+1})\|^{2}-\|\mathbf{G}(\mathbf{y}^{k})\|^{2})$
$+\frac{\tilde{\beta}^{2}}{2}\|\mathbf{y}^{k+1}-\mathbf{y}^{k}\|^{2}.$
Substituting the above two inequalities into (45) gives
$\mathbf{f}(\mathbf{y}^{k+1})-\mathbf{f}(\mathbf{y}^{\star})\leq S^{k}-S^{k+1}$
$+\frac{1}{2}(\tilde{\beta}^{2}+L_{\mathcal{R}^{k}}-\frac{1}{\gamma})\|\mathbf{y}^{k+1}-\mathbf{y}^{k}\|^{2}.$ (46)
Next, note from Lemma 2 that
$L_{\mathcal{R}^{k}}\leq L_{F}+\frac{1}{\rho}\|\mathbf{B}^{T}\mathbf{B}\|+\sqrt{np}(\|\mathbf{q}^{k}\|+\|\mathbf{G}(\mathbf{y}^{k})\|)L_{g}$
$\leq L_{F}+\frac{1}{\rho}\|\mathbf{B}^{T}\mathbf{B}\|+2\sqrt{np}\max\{\|\mathbf{q}^{k}\|,\|\mathbf{G}(\mathbf{y}^{k})\|\}L_{g}.$
Due to (41), $\|\mathbf{G}(\mathbf{y}^{0})\|\leq 2(\|\bm{\lambda}^{\star}\|+\sqrt{C})$, and the hypothesis of Lemma 4, we have
$L_{\mathcal{R}^{k}}\leq L_{F}+\frac{1}{\rho}\|\mathbf{B}^{T}\mathbf{B}\|+4\sqrt{np}(\|\bm{\lambda}^{\star}\|+\sqrt{C})L_{g},$
for each $k=0,1,\ldots,K-1$. This, along with Lemma 2 and the parameter condition in Theorem 1, gives $\frac{1}{\gamma}\geq\tilde{\beta}^{2}+L_{\mathcal{R}^{k}}$ $\forall k=0,\ldots,K-1$. It then follows from (46) that $\mathbf{f}(\mathbf{y}^{k+1})-\mathbf{f}(\mathbf{y}^{\star})\leq S^{k}-S^{k+1}$ $\forall k=0,1,\ldots,K-1$, leading to (42). ∎
Based on the above lemmas, we provide the bounds on $\|\mathbf{q}^{k}\|$ and $\|\mathbf{u}^{k}-\mathbf{u}^{0}\|_{W}$ using mathematical induction.
Lemma 5.
Suppose all conditions in Theorem 1 hold. For each $k\geq 0$,
$\|\mathbf{q}^{k}\|\leq 2(\|\bm{\lambda}^{\star}\|+\sqrt{C}),$ (47)
$\|\mathbf{u}^{k}-\mathbf{u}^{0}\|_{W}\leq\|\mathbf{u}^{0}\|_{W}+\|\mathbf{u}^{\star}\|_{W}+\sqrt{\frac{2C}{\rho}}.$ (48)
Proof.
We first prove (47) for all $k\geq 0$ by induction. Due to (41), $\|\mathbf{q}^{0}\|\leq\|\mathbf{G}(\mathbf{y}^{0})\|\leq 2\sqrt{C}$, so that (47) is satisfied when $k=0$. Next, given $K\geq 1$, suppose (47) holds for all $k=0,1,\ldots,K-1$, and below we show (47) holds for $k=K$.
To do so, note from Proposition 1(e) that $\mathbf{f}(\mathbf{y}^{\star})=\operatorname{\min}_{\mathbf{y}\in Y}~{}\mathbf{f}(\mathbf{y})+\langle\bm{\lambda}^{\star},\mathbf{G}(\mathbf{y})\rangle+\langle\mathbf{u}^{\star},\mathbf{B}\mathbf{y}-\mathbf{c}\rangle$, implying that $\mathbf{f}(\mathbf{y}^{\star})\leq\mathbf{f}(\mathbf{y}^{k})+\langle\bm{\lambda}^{\star},\mathbf{G}(\mathbf{y}^{k})\rangle+\langle\mathbf{u}^{\star},\mathbf{B}\mathbf{y}^{k}-\mathbf{c}\rangle$ $\forall k\geq 1$. Adding this inequality over $k=1,\ldots,K$ yields
$\sum_{k=1}^{K}(\mathbf{f}(\mathbf{y}^{\star})-\mathbf{f}(\mathbf{y}^{k}))$
$\leq\langle\bm{\lambda}^{\star},\sum_{k=1}^{K}\mathbf{G}(\mathbf{y}^{k}\rangle+\langle\mathbf{u}^{\star},\sum_{k=1}^{K}(\mathbf{B}\mathbf{y}^{k}-\mathbf{c})\rangle.$ (49)
By referring to the proof of [25, eq.(40), eq.(41)], we can show that $\sum_{k=1}^{K}\mathbf{G}(\mathbf{y}^{k})\leq\mathbf{q}^{K}$ and $\sum_{k=1}^{K}(\mathbf{1}_{n}\otimes I_{m+p})^{T}(\mathbf{B}\mathbf{y}^{k}-\mathbf{c})=\rho(\mathbf{1}_{n}\otimes I_{m+p})^{T}(\mathbf{u}^{K}-\mathbf{u}^{0})$. Also note that $\bm{\lambda}^{\star}\geq\mathbf{0}$ and $\mathbf{u}^{\star}\in\operatorname{Range}(\mathbf{1}_{n}\otimes I_{m+p})$. It follows from (49) that
$\sum_{k=1}^{K}(\mathbf{f}(\mathbf{y}^{\star})-\mathbf{f}(\mathbf{y}^{k}))\leq\langle\bm{\lambda}^{\star},\mathbf{q}^{K}\rangle+\rho\langle\mathbf{u}^{\star},\mathbf{u}^{K}-\mathbf{u}^{0}\rangle.$
By combining the above inequality with (42) and by using $W\mathbf{u}^{\star}=\mathbf{u}^{\star}$ and $\langle W\mathbf{u}^{\star},\mathbf{u}^{K}-\mathbf{u}^{0}\rangle=\langle W^{\frac{1}{2}}\mathbf{u}^{\star},W^{\frac{1}{2}}(\mathbf{u}^{K}-\mathbf{u}^{0})\rangle$,
$S^{K}\leq S^{0}\!+\!\langle\bm{\lambda}^{\star},\mathbf{q}^{K}\rangle+\rho\langle W\mathbf{u}^{\star},\mathbf{u}^{K}-\mathbf{u}^{0}\rangle$
$\leq S^{0}+\|\bm{\lambda}^{\star}\|\cdot\|\mathbf{q}^{K}\|+\rho\|\mathbf{u}^{\star}\|_{W}(\|\mathbf{u}^{K}\|_{W}+\|\mathbf{u}^{0}\|_{W}).$ (50)
In addition, following the proof of [25, Theorem 1], we can show via Proposition 1(c) that $\|\mathbf{G}(\mathbf{y}^{K})\|^{2}\leq\frac{1}{2}\|\mathbf{q}^{K}\|^{2}+\tilde{\beta}^{2}\|\mathbf{y}^{K}-\mathbf{y}^{\star}\|^{2}+\|\mathbf{G}(\mathbf{y}^{\star})\|^{2}$. By incorporating this into the expression of $S^{K}$ given in Lemma 4 and because $\frac{1}{\gamma}\geq\tilde{\beta}^{2}+\frac{1}{\rho}\|\mathbf{B}^{T}\mathbf{B}\|$,
$S^{K}\geq\frac{\rho}{2}\|\mathbf{u}^{K}\|_{W}^{2}+\frac{1}{4}\|\mathbf{q}^{K}\|^{2}-\frac{1}{2}\|\mathbf{G}(\mathbf{y}^{\star})\|^{2}.$ (51)
Thus, (50) and (51) together result in
$\frac{\rho}{2}(\|\mathbf{u}^{K}\|_{W}-\|\mathbf{u}^{\star}\|_{W})^{2}+\frac{1}{4}(\|\mathbf{q}^{K}\|-2\|\bm{\lambda}^{\star}\|)^{2}\leq C.$ (52)
It can be seen that $\|\mathbf{q}^{K}\|\leq 2(\|\bm{\lambda}^{\star}\|+\sqrt{C})$. This completes the proof of (47) for each $k\geq 0$.
Finally, note from (52) that $\|\mathbf{u}^{K}-\mathbf{u}^{0}\|_{W}\leq\|\mathbf{u}^{0}\|_{W}+\|\mathbf{u}^{\star}\|_{W}+(\|\mathbf{u}^{K}\|_{W}-\|\mathbf{u}^{\star}\|_{W})\leq\|\mathbf{u}^{0}\|_{W}+\|\mathbf{u}^{\star}\|_{W}+\sqrt{\frac{2C}{\rho}}$. Since $K$ can be any positive integer and (48) is satisfied when $k=0$, we conclude that (48) holds for all $k\geq 0$. ∎
From (47), $\mathbf{q}^{k}\leq 2(\|\bm{\lambda}^{\star}\|+\sqrt{C})\mathbf{1}_{np}$ $\forall k\geq 0$, which, together with (37), leads to (25). In addition, since $P^{W}=(P^{W})^{T}\succeq\mathbf{O}$ and $P^{W}\mathbf{1}_{n}=\mathbf{1}_{n}$, we have $P^{W}-\frac{\mathbf{1}_{n}\mathbf{1}_{n}^{T}}{n}\succeq\mathbf{O}$. Hence, $\|(\mathbf{1}_{n}\otimes I_{m+p})^{T}(\mathbf{u}^{k}-\mathbf{u}^{0})\|^{2}\leq n(\mathbf{u}^{k}-\mathbf{u}^{0})^{T}(P^{W}\otimes I_{m+p})(\mathbf{u}^{k}-\mathbf{u}^{0})=n\|\mathbf{u}^{k}-\mathbf{u}^{0}\|^{2}_{W}$. It follows from (38) and (48) that (26) holds.
-C Proof of Theorem 2
Let $k\geq 1$. Because $\frac{I}{\gamma}-\frac{\mathbf{B^{T}}\mathbf{B}}{\rho}\succ\mathbf{O}$ and because of (41), we have $S^{k}\geq 0$, where $S^{k}$ is defined in Lemma 4. This, along with (42), results in $\sum_{t=1}^{k}(\mathbf{f}(\mathbf{y}^{t})-\mathbf{f}(\mathbf{y}^{\star}))\leq S^{0}$. Moreover, due to the convexity of $\mathbf{f}$, $\mathbf{f}(\bar{\mathbf{y}}^{k})-\mathbf{f}(\mathbf{y}^{\star})\leq\frac{1}{k}\sum_{t=1}^{k}(\mathbf{f}(\mathbf{y}^{t})-\mathbf{f}(\mathbf{y}^{\star}))$. Therefore, $\mathbf{f}(\bar{\mathbf{y}}^{k})-\mathbf{f}(\mathbf{y}^{\star})\leq\frac{S^{0}}{k}$.
On the other hand, since $\mathbf{f}(\mathbf{y}^{\star})=\min_{\mathbf{y}\in Y}~{}\mathbf{f}(\mathbf{y})+\langle\bm{\lambda}^{\star},\mathbf{G}(\mathbf{y})\rangle+\langle u^{\star},(\mathbf{1}_{n}\otimes I_{m+p})^{T}(\mathbf{B}\mathbf{y}-\mathbf{c})\rangle$,
$\mathbf{f}(\bar{\mathbf{y}}^{k})-\mathbf{f}(\mathbf{y}^{\star})\geq$ $-\langle\bm{\lambda}^{\star},\mathbf{G}(\bar{\mathbf{y}}^{k})\rangle$
$-\langle u^{\star},(\mathbf{1}_{n}\otimes I_{m+p})^{T}(\mathbf{B}\bar{\mathbf{y}}^{k}-\mathbf{c})\rangle.$ (53)
Because $\bm{\lambda}^{\star}\geq\mathbf{0}$ and because of (25),
$\langle\bm{\lambda}^{\star},\mathbf{G}(\bar{\mathbf{y}}^{k})\rangle\leq\frac{2(\|\bm{\lambda}^{\star}\|+\sqrt{C})\mathbf{1}_{np}^{T}\bm{\lambda}^{\star}}{k}.$ (54)
Additionally, since $\mathbf{u}^{\star}=\mathbf{1}_{n}\otimes u^{\star}$ and due to (26),
$\langle u^{\star},(\mathbf{1}_{n}\otimes I_{m+p})^{T}(\mathbf{B}\bar{\mathbf{y}}^{k}-\mathbf{c})\rangle$
$\leq$ $\|u^{\star}\|\cdot\|(\mathbf{1}_{n}\otimes I_{m+p})^{T}(\mathbf{B}\bar{\mathbf{y}}^{k}-\mathbf{c})\|$
$\leq$ $\frac{\rho\|\mathbf{u}^{\star}\|(\|\mathbf{u}^{0}\|_{W}+\|\mathbf{u}^{\star}\|_{W}+\sqrt{2C/\rho})}{k}.$ (55)
Incorporating (54) and (55) into (53) yields $\mathbf{f}(\bar{\mathbf{y}}^{k})-\mathbf{f}(\mathbf{y}^{\star})\geq-\frac{C^{0}}{k}$. This completes the proof.
-D Proof of Corollary 1
Let $k\geq 1$. From the definitions of $\mathbf{B}$ and $\mathbf{c}$ in Section III,
$\|(\mathbf{1}_{n}\otimes I_{m+p})^{T}(\mathbf{B}\bar{\mathbf{y}}^{k}-\mathbf{c})\|^{2}$
$=$ $\|\sum_{i\in\mathcal{V}}\bar{t}_{i}^{k}\|^{2}+\|\sum_{i\in\mathcal{V}}\Bigl{(}\bigl{(}\sum_{j\in\mathcal{N}_{i}}A_{ji}\bar{x}_{i}^{k}\bigr{)}-b_{i}\Bigr{)}\|^{2}$
$=$ $\|\sum_{i\in\mathcal{V}}\bar{t}_{i}^{k}\|^{2}+\|\sum_{i\in\mathcal{V}}(A_{i}\bar{x}_{\mathcal{N}_{i}}^{k}-b_{i})\|^{2},$ (56)
where $\bar{t}_{i}^{k}=\frac{1}{k}\sum_{\ell=1}^{k}t_{i}^{\ell}$. This implies $\|\sum_{i\in\mathcal{V}}(A_{i}\bar{x}_{\mathcal{N}_{i}}^{k}-b_{i})\|\leq\|(\mathbf{1}_{n}\otimes I_{m+p})^{T}(\mathbf{B}\bar{\mathbf{y}}^{k}-\mathbf{c})\|$, which, together with (26), yields (29). In addition, (56) also leads to $\|\sum_{i\in\mathcal{V}}\bar{t}_{i}^{k}\|\leq\|(\mathbf{1}_{n}\otimes I_{m+p})^{T}(\mathbf{B}\bar{\mathbf{y}}^{k}-\mathbf{c})\|$. By combining this, (25), and (26) with $\sum_{i\in\mathcal{V}}g_{i}(\bar{x}^{k}_{\mathcal{N}_{i}})\leq\sum_{i\in\mathcal{V}}(g_{i}(\bar{x}^{k}_{\mathcal{N}_{i}})-\bar{t}_{i}^{k})+\|\sum_{i\in\mathcal{V}}\bar{t}_{i}^{k}\|\mathbf{1}_{p}$, we obtain (28). Finally, (30) can be directly derived from (27).